### v1

# Recursions, Trains, Trees, and Combinatorial Rod Set Algebra

Abstract
We explore a physical model of ordered sums of integers as trains of rods. The trains for a fixed, possibly infinite, set of rod lengths naturally correspond to nodes in a tree; relations among finite linear recursions encoded in the subtrees define algebraic operations on sets of rods. We use this algebra to prove classic identities for recursively defined sequences, to show that Lucas sequences are divisibility sequences, to characterize two-term linear Fibonacci identities, and to find the cyclotomic polynomial factors of Borwein trinomials. We complement abstractions with lots of examples.
Introduction
Children use Cuisenaire™ rods[3] to learn how numbers fit together. A common activity is to make trains consisting of a sequence of rods placed end-to-end, and to count how many trains are possible under various restrictions. Figure 1 shows all possible trains of length $10$ made from rods of lengths $2$, 33, and $5$. The trains model compositions: sequences of positive integers with a given sum.
Figure 1: The $14$ trains of length $10$ made from Cuisenaire™ rods of lengths 22 (red), 33 (green), and 55 (yellow). Each row is a separate train.
The Fibonacci numbers count the ways to build a train of length $n$ from rods of lengths $1$ and $2$; the recursion
$\operatorname{F}(n)=\operatorname{F}(n-1)+\operatorname{F}(n-2)$
sorts the trains of length $n$ depending on the length of the first (or last) rod. In this model for the Fibonacci sequence, indexing starts at $1$ with $\operatorname{F}(1)=1$ and $\operatorname{F}(2)=2$.
Playing with the recursion, we see that
$\operatorname{F}(n)$ $=\operatorname{F}(n-1)+\operatorname{F}(n-2)$
$=\operatorname{F}(n-1)+(\operatorname{F}(n-2-1)+\operatorname{F}(n-2-2))$
$=\operatorname{F}(n-1)+\operatorname{F}(n-3)+\operatorname{F}(n-4).$
It follows that for $n>2$, the sequence $\operatorname{F}(n)$ also satisfies the recursion determined by rods of length $1$, 33, and $4$.
Soon we will use our rod model to study this observation about the Fibonacci sequence by converting trains composed of rods of lengths $1$ and $2$ into trains composed of rods of lengths $1$, 33, and $4$. When a 22 is followed by a $1$ we can glue them together to make a $3$, and when 22 is followed by a $2$ we glue to make a $4$. This construction may end with an extra rod of length $2$ at the end of the train. Figure 2 shows an example.
Figure 2: Converting the top train of $1$’s and 22’s into the bottom train of 11s, 33s, and 44s, with a 22 at the end.
Kids can’t play with infinitely many rods, but we can. If we repeatedly replace the rods of even length $2k$ by rods of lengths $2k+1$ and $2k+2$, we see that when $n$ is odd, $\operatorname{F}(n)$ satisfies the recursion defined by rods of length $1,3,5,7,\ldots]$.
Kids can identify Cuisenaire™ rods by color or by length. But writing
$\operatorname{F}(n)$ $=\operatorname{F}(n-1)+\operatorname{F}(n-2)$
$=(\operatorname{F}(n-1-1)+\operatorname{F}(n-1-2))+\operatorname{F}(n-2)$
$=2\operatorname{F}(n-2)+\operatorname{F}(n-3),$
shows that when $n>3$, the Fibonacci train counts also satisfy the recursion for trains built from two different kinds of rods of length $2$ along with rods of length $3$. To preserve our physical model we will allow rods with the same length but different colors.
Since
$\operatorname{F}(n-1)=\operatorname{F}(n-2)+\operatorname{F}(n-3),$
we know that
$\operatorname{F}(n-2)=\operatorname{F}(n-1)-\operatorname{F}(n-3),$
so
$\operatorname{F}(n)$ $=\operatorname{F}(n-1)+\operatorname{F}(n-2)$
$=\operatorname{F}(n-1)+(\operatorname{F}(n-1)-\operatorname{F}(n-3))$
$=2\operatorname{F}(n-1)-\operatorname{F}(n-3),$
when $n>3$, which suggests that we will also need a way to define “negative” rods (which will not be rods of negative length).
In the first five sections we’ll formalize this rod train model for relationships between recursions. Section 1 has careful definitions for rods and rod trains. Sections 2 and 3 introduce a tree structure for rod trains that leads to a natural way to understand the recursion manipulations in this Introduction. In Sections 4 and 5 we introduce algebraic operations on rod sets.
In Section 6 we pause to connect our work to generating functions; we see there that rod set algebra provides a combinatorial interpretation for polynomial multiplication. In Section 7 we define duality, the last operation on rod sets, and use it to prove many identities. The next six sections cover applications, often involving small finite rod sets. We study periodic sequences and connect patterns of train counts to rod set operations, which leads to more Fibonacci identities, our proof that Lucas sequences are divisibility sequences, and rediscovering the cyclotomic factors of Borwein trinomials. We end with further avenues to explore.
1 Setting the Stage
We start with careful definitions and some elementary observations about rods, rod sets and rod trains.
Definition 1.
A rod is a triple $(r,c,\pm 1)$ in which the length $r$ is a positive integer and $c$ is an arbitrary tag we think of as a color. The third attribute is the sign of the rod. A rod of length $r$ with negative sign is an antirod, written as $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{r}$}. Sometimes “rod” will mean a rod of either sign and sometimes just a rod with positive sign. The context should make clear when the sign is important.
Often we use the length of a rod $r$ in an arithmetic expression such as $n-\operatorname{length}(r)$. When no confusion can result we will omit the $\operatorname{length}$ function and write simply nrn-r. Formally, if $r=(m,c,\pm 1)$ then nr=nmn-r=n-m.
Definition 2.
A rod train, or simply a train, is a finite sequence of rods. The empty train is the empty sequence $\Lambda$. For the nonempty train $\tau$ we write $|\tau|$ for the rod with length the sum of the lengths of the rods in $\tau$, sign the product of the signs, and color that distinguishes it from other rods in its context. We can think of this operation as gluing together the rods in $\tau$. Since by definition rods have positive length there is no rod $|\Lambda|$.
The length and sign of train $\tau$ are the length and sign of rod $|\tau|$. Thus a train is positive when it contains an even number of antirods and negative otherwise.
We concatenate trains $\sigma$ and $\tau$ as either $\sigma\tau$ or $\sigma.\tau$; we use the dot when we want it for emphasis. In general $\sigma\tau\neq\tau\sigma$, but $|\sigma\tau|=|\tau\sigma|$ always.
For example, $|{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}}|=|1{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}1|=|11{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}|=|1.\Lambda.1{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}|={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}}$.
Definition 3.
A rod set is a set of rods with finitely many rods of each length. We write a rod set as a square bracketed list of rods, usually in non-decreasing length order. Duplicate entries implicitly specify rods of different colors. We write $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{R}$} for the rod set whose rods reverse the signs of the rods in $R$.
The concatenation of rod sets $R$ and $S$ is the rod set
$RS=SR=\{|jk|\mid j\in R,k\in S\}.$
Let #R\#R represent the number of rods in a finite rod set $R$. Let $R^{+}(n)$ and $R^{-}(n)$ be the sets of positive and negative rods of length $n$ in $R$. The net rod count $\operatorname{C}(n,R)$ for $n>0$ is:
$\operatorname{C}(n,R)=\#{R^{+}(n)}-\#R^{-}(n).$
The net rod count $C(n,R)$ can be positive, negative, or zero. Since there are no rods with length 0 we could consistently define $\operatorname{C}(0,R)=0$. For reasons that will become clear later on, we choose instead to leave it undefined.
When there are multiple rods of the same length we may use exponents to count them, so
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}},2,2,2]=[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}}^{2},2^{3}].$
In rod set unions we assume that colors are all distinct, so
$2,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}},3]\cup[2,2,4]=[2^{3},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}},3,4].$
In intersections and set differences we assume colors match as much as possible, so
$2,2,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}},3]\cap[2,2,4]=[2,2]$
and
$2,2,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}},3]\setminus[2]=[2,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}},3].$
Definition 4.
A rod train set, or simply a train set, is a set of rod trains. The net train count $\operatorname{C}(n,T)$ is the difference between the number of positive and negative trains of that length.
For train set $T$ we write
$|T|=\{|\tau|\mid\tau\in T\}$
for the set of rods built from the nonempty trains in $T$. Thus $\operatorname{C}(n,T)=\operatorname{C}(n,|T|)$ when $n>0$.
The set of trains built from the rods in rod set $R$ is $\operatorname{Trains}(R)$, and $|\operatorname{Trains}(R)|$ is the rod set whose rods are built from the following trains:
$\operatorname{Trains}(R)=\{\Lambda\}\cup R\cup R.R\cup R.R.R\cup\cdots.$ (1.1)
$|\operatorname{Trains}([\phantom{a}])|$ is empty because $\operatorname{Trains}([\phantom{a}])=\{\Lambda\}$ and there is no rod $|\Lambda|$.
Definition 5.
For rod set $R$, the net train count sequence is
$\operatorname{F}(n,R)=\operatorname{C}(n,\operatorname{Trains}(R))=\begin{cases}\operatorname{C}(n,|\operatorname{Trains}(R)|)&n>0\\ 1&n=0\\ 0&n<0.\end{cases}$
We usually think of this sequence as starting at $n=0$. The 0 values for $n<0$ are there to make some algebra cleaner.
When the rod set $R$ is clear from context we may shorten expressions by writing $\operatorname{F}(n)$ instead of $\operatorname{F}(n,R)$.
The traditional function identifier, $\operatorname{F}$,” for the function that counts trains is a nice nod to the start of our study. It’s short for Fibonacci: the sequence of train counts for rods of lengths $1$ and $2$.
Table 1 displays the positive and negative trains of length at most $5$ for the rod set $1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}]$. Ignoring signs, the total number of trains of each length is, of course, the corresponding Fibonacci number. The last column is the net train count.
Length
Positive
Negative
Train
Count
$\operatorname{F}(n,[1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}])$
0 $\Lambda$ 1 1
1 1 1 1
2 11 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}$} 2 0
[…]
Table 1: Rod trains for $1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}]$
Theorem 6.
For rod set $R$, the net train count sequence $\operatorname{F}(n,R)$ satisfies the recursion
$\operatorname{F}(n,R)=\sum_{k\in R}(\operatorname{sign}k)\operatorname{F}(n-k,R),$ (1.2)
with initial conditions
$\operatorname{F}(n,R)=\begin{cases}1\text{ if }n=0\\ 0\text{ if }n<0\ .\end{cases}$ (1.3)
Proof.
Let $\operatorname{F}^{+}(n,R)$ be the number of positive trains in $\operatorname{Trains}(n,R)$ and $\operatorname{F}^{-}(n,R)$ the number of negative trains. Then
$\operatorname{F}(n,R)=\operatorname{F}^{+}(n,R)-\operatorname{F}^{-}(n,R).$
Let $k$ be the length of the last rod in a train $\tau$ of length $n>0$. Removing that rod from $\tau$ leaves a train $\tau^{\prime}$ of length nkn-k. If $k$ is an antirod the signs of $\tau$ and $\tau^{\prime}$ are opposite, otherwise they are the same.
Let $R^{+}$ and $R^{-}$ be the sets of rods and antirods in $R$. Then
$F^{+}(n,R)=\sum_{k\in R^{+}}F^{+}(n-k,R)+\sum_{k\in R^{-}}F^{-}(n-k,R).$
and
$F^{-}(n,R)=\sum_{k\in R^{-}}F^{+}(n-k,R)+\sum_{k\in R^{+}}F^{-}(n-k,R).$
Note that if $k>n$ the initial conditions for negative $n$ ensure that no extra terms are added.
Thus
$\operatorname{F}(n,R)$ $=\operatorname{F}^{+}(n,R)-\operatorname{F}^{-}(n,R)$
$=\sum_{k\in R^{+}}\Bigl{(}\operatorname{F}^{+}(n-k,R)-\operatorname{F}^{-}(n-k,R)\Bigr{)}$
$\ \ \ \ \ -\sum_{k\in R^{-}}\Bigl{(}\operatorname{F}^{+}(n-k,R)-\operatorname{F}^{-}(n-k,R)\Bigr{)}$ (1.4)
$=\sum_{k\in R}\Bigl{(}(\operatorname{sign}k)(\operatorname{F}^{+}(n-k,R)-\operatorname{F}^{-}(n-k,R)\Bigr{)}.$
The initial condition $\operatorname{F}(0,R)=1$ guarantees that when $k\in R$ the one rod train $k$ contributes $\pm 1$ to the net train count. It is consistent with the fact that the empty train contains an even number of antirods. At the end of Section 5 we show how to use rod trains to model recursive sequences with different initial conditions.
In Benjamin and Quinn’s Proofs that Really Count[2], rod sets with these default initial conditions correspond to weighted unphased tilings of a board. Because we allow infinite rod sets and are particularly interested in the relations among rod sets, we focus on the rods themselves rather than introducing a board for them to tile. Their system of weights accommodates any complex number, whereas we only need weights of $\pm 1$. Hopkins [8] uses Cuisenaire rods as we do, with many engaging classroom activities, but without antirods or the connections among rod sets.
Here are some familiar sequences described as net train counts, starting at $n=0$:
  • $\operatorname{F}(n,[1,2])=1,1,2,3,5,8,\ldots$, the Fibonacci numbers.
  • $\operatorname{F}(n,[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}},2])=1,-1,2,-3,5,-8,\ldots$, the alternating Fibonacci numbers.
  • $\operatorname{F}(n,[1,3,5,\ldots])=1,1,1,2,3,5,8,\ldots$, the Fibonacci sequence with an extra initial $1$.
  • $\operatorname{F}(n,[1])=1,1,1,1,\ldots$, the constant sequence of all $1$’s.
  • $\operatorname{F}(n,[1,1])=1,1,2,4,8,\ldots$, the powers of $2$.
  • $\operatorname{F}(n,[2,3])=1,0,1,1,1,2,2,3,4,5,7,9,12,\ldots$, a Padovan sequence [14].
  • $\operatorname{F}(n,[1,3])=1,1,1,1,2,3,4,6,9,13,19,\ldots$, a Narayana cow number sequence [13].
  • $\operatorname{F}(n,[1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}])=1,1,0,-1,-1,0,1,\ldots$, repeating with period $6$.
  • $\operatorname{F}(n,[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}])=1,1,0,1,1,0,1,\ldots$, repeating with period $3$.
  • $\operatorname{F}(n,[\phantom{a}])=\operatorname{F}(n,[1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}}])=1,0,0,0,\ldots$ .
The next theorem shows that studying the net train count sequences of rod sets is quite general.
Theorem 7.
For every infinite sequence of integers $a_{1},a_{2},\ldots$ there is a rod set $R$ such that $\operatorname{F}(n,R)=a_{n}$ for all $n\geq 1$.
Proof.
We construct $R$ inductively. By definition, $\operatorname{F}(0,R)=1$. To make $\operatorname{F}(1,R)=a_{1}$, the set $R$ clearly needs $a_{1}$ rods of length $1$ (these will be antirods if $a_{1}<0$). Once the number of rods of length $k$ is known for all $k<n$, we can use the recursion determined by these rods to compute the number of trains of length $n$ composed of more than one rod. Then we add the correct number of rods or antirods of length $n$ so that the new one-rod trains make $\operatorname{F}(n,R)=a_{n}$. ∎
For the train counts corresponding to the Catalan numbers,
$\operatorname{F}(n,R)=1,1,2,5,14,42,132,429,\ldots,$
the corresponding rod set is
$R=[1,2,3^{2},4^{5},5^{14},6^{42},7^{132},8^{429},\ldots].$
The number of rods of each length is the Catalan number with index shifted by $1$ from the train counts. The recursion corresponds to the well-known convolution that the Catalan’s satisfy.
2 Trees
Here we explore a natural way to organize $\operatorname{Trains}(R)$ in a tree111 Sorry for the awkward metaphor mixing trees and trains. that lets us sytematically study the way we combined rods in Figure 2.
Definition 8.
Let $R$ be a rod set. $\operatorname{Tree}(R)$ is the rooted tree in which each node has one child for each rod in $R$. We label the edge from parent to child by that rod. Then we label each node by the rod train that is the path from the root to that node. The root’s label is the empty train $\Lambda$. The children of a node labeled by the train $\tau$ are labeled by the trains $\tau.R$ constructed by appending a rod $k\in R$ to $\tau$. The nodes at the $m$th level correspond to trains containing $m$ rods.
When $\tau\in\operatorname{Trains}(R)$ and no confusion will result, we may write “node $\tau$” for the corresponding node in $\operatorname{Tree}(R)$ instead of the more formal “node with label $\tau$.”
We decorate node $\tau$ by adding $|\tau|$ as a second label. Since gluing trains of the same length together yields rods of the same length, these second labels do not identify nodes.
In the language of theoretical computer science, $R$ is an alphabet and $\operatorname{Trains}(R)$ are the words or strings in the language for that alphabet. The tree structure depends only on the number of letters in the alphabet (which may be infinite). When $R$ contains just two rods it’s the complete binary tree.
Figure 3 shows the first four levels of $\operatorname{Tree}([1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}])$. The Fibonacci numbers count the number of node labels of each length. We can use this figure to confirm the entries in Table 1.
Figure 3: The first four levels of $\operatorname{Tree}([1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}])$
Figure 2 in the Introduction corresponds to replacing the node $\tau=2$ at the first level of $\operatorname{Tree}([1,2])$ by its children $21$ and $22$. We write this operation on nodes as
$\{1,2\}\xrightarrow{\hphantom{a}2\hphantom{a}}\{1,21,22\}.$ (2.1)
We use Equation 2.1 to replace some pairs of rods in an element of $\operatorname{Trains}([1,2])$ with an element of $\operatorname{Trains}([1,3,4])$, possibly with an extra rod of length $2$ at the end. For the example in Figure 2 the parsing is
$\tau=12221211122221212=1.22.21.21.1.1.22.22.1.21.2,$ (2.2)
which maps $\tau$ to the element $1433114413$ of $\operatorname{Trains}([1,3,4])$ with an appended rod of length $2$.
We can iterate this kind of tree node replacement. For example, replacing node $1$ next and then node $21$ leads to the expansion chain we write as
$\{1,2\}\xrightarrow{\hphantom{a}2\hphantom{a}}\{1,21,22\}\xrightarrow{\hphantom{a}1\hphantom{a}}\{11,12,21,22\}\xrightarrow{\hphantom{a}21\hphantom{a}}\{11,12,211,212,22\}.$ (2.3)
The next definition formalizes this kind of expansion.
Definition 9.
A rooted subtree $\mathfrak{T}$ of $\operatorname{Tree}(R)$ is an expansion subtree when
  • It contains all the children of the root.
  • If it contains a child of some node, then it contains all the children of that node.
For an expansion subtree $\mathfrak{T}$, $\text{Inner}(\mathfrak{T})$ is the set of nodes (other than the root) of $\mathfrak{T}$ that have at least one child (hence all children) in $\mathfrak{T}$; we call these the internal nodes. $\text{Leaves}(\mathfrak{T})$ is the set of nodes of $\mathfrak{T}$ that have no children in $\mathfrak{T}$.
We write the expansion corresponding to expansion subtree $\mathfrak{T}$ as
$R\xrightarrow{\hphantom{a}\text{Inner}(\mathfrak{T})\hphantom{a}}\text{Leaves}(\mathfrak{T}).$
The expansion subtree $\mathfrak{T}$ whose only nodes are the children of the root has $\text{Leaves}(\mathfrak{T})=R$. If every node of $\operatorname{Tree}(R)$ is internal in $\mathfrak{T}$ then $\text{Leaves}(\mathfrak{T})$ is empty.
Figure 4 shows three expansions; in each figure the internal nodes are circled and the leaves boxed:
$\{1,2\}$ $\xrightarrow{\hphantom{a}\{1,2,21\}\hphantom{a}}\{11,12,22,211,212\},$
$\{1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}\}$ $\xrightarrow{\hphantom{a}\{1,1{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}},1{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}1\}\hphantom{a}}\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}},11,1{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}},1{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}11,1{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}1{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}\},$
$\{1,2\}$ $\xrightarrow{\hphantom{a}\{2,22,222,\ldots\}\hphantom{a}}\{1,21,221,2221,\ldots\}.$
$a$ Refer to caption $a$ Refer to caption
Figure 4: Expansions of $[1,2]$ and $1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}]$, with internal nodes circled, leaves boxed.
3 Discrepancies
In the Introduction we saw that the Fibonacci train counts $\operatorname{F}(n,[1,2])$ satisfy the recursion for the rod set $[1,3,4]$ except when $n=2$, which was the single rod we replaced in Expansion 2.1.
In this section we show that the internal nodes in an expansion subtree correspond to the places where the original net train counts fail to satisfy the recursion determined by the expansion.
We begin with a definition that refers only to rod sets and recursions, not to subtrees.
Definition 10.
For rod sets $R$ and $S$ the discrepancy at $n>0$ is the difference
$\operatorname{D}(n,R,S)=\operatorname{F}(n,R)-\sum_{k\in S}(\operatorname{sign}k)\operatorname{F}(n-k,R).$
Thus $\operatorname{D}(n,R,S)$ measures by how much the net train counts for $R$ fail to satisfy the recursion for $S$ at $n$. We look at some examples to get a feel for discrepancies and to motivate the theorem that will characterize them.
First, $\operatorname{D}(n,[1,2],[1,3,4])$ is $1$ when $n=2$ and is 0 everywhere else.
Clearly $\operatorname{D}(n,R,R)=0$ for all $n>0$.
Earlier we asserted without proof that the net train counts
$\operatorname{F}(n,[1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}])=1,1,0,-1,-1,0,\ldots$
repeat with period $6$. A sequence with period $6$ satisfies the recursion defined by the rod set $[6]$, so to prove that assertion we need to show that for $n>0$ the discrepancies satisfy
$\operatorname{D}(n,[1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}],[6])=\begin{cases}1&n=1\\ -1&n=3,4\\ 0&\text{otherwise}.\end{cases}$ (3.1)
This will follow from the next theorem, which will also imply the following two Fibonacci identities. The sequence $\operatorname{F}(n)=\operatorname{F}(n,[1,2])=1,1,2,3,5,8,13,\ldots$, written here with our default initial conditions, shifts the indices of the traditional initial conditions by one. With this notation
$1+2+5+13+34+\ \cdots\ +\operatorname{F}(2n)$ $=\operatorname{F}(2n+1),$ (3.2)
$1+3+8+21+\cdots+\operatorname{F}(2n-1)$ $=\operatorname{F}(2n)-1.$ (3.3)
Since the recursion for the rod set $\text{odds}=[1,3,5,\ldots]$ is
$\operatorname{F}(n,\text{odds})=\operatorname{F}(n-1,\text{odds})+\operatorname{F}(n-3,\text{odds})+\cdots$
the above identities are equivalent to
$\operatorname{D}(n,[1,2],\text{odds})=\begin{cases}0&\text{if $n$ is odd}\\ 1&\text{if $n$ is even},\end{cases}$
which will follow from the next theorem.
Lemma 11.
(Parsing) Suppose $\mathfrak{T}$ is an expansion subtree of $\operatorname{Tree}(R)$. Then we can write every nonempty $R$-train $\rho$ uniquely as a concatenation
$\rho=\sigma_{1}.\sigma_{2}.\cdots.\sigma_{m}.\tau$
where each $\sigma_{i}\in\text{Leaves}(\mathfrak{T})$ and $\tau\in\text{Inner}(\mathfrak{T})$. In this expression $m$ may be 0 (that is, no $\sigma_{i}$s) and $\tau$ may be $\Lambda$, but not both.
Conversely, every such expression is an $R$-train.
Proof.
The argument generalizes the example in Equation 2.2. Given $\rho\in\operatorname{Trains}(R)$, we follow $\rho$ from the root of $\operatorname{Tree}(R)$ until we encounter a node $\sigma_{1}\in\text{Leaves}(\mathfrak{T})$. We continue inductively with the rest of $\rho$ to find $\sigma_{2},\ldots\sigma_{m}$. The construction ends when there are no rods left, or too few left to reach a node in $\text{Leaves}(\mathfrak{T})$. Then the trailing train ends at a node $\tau\in\text{Inner}(\mathfrak{T})$.
The converse is the simple observation that each $\sigma_{i}$ is just a sequence of rods from $R$. ∎
Using the second expansion tree in Figure 4 we can see the parsings
[…]
Theorem 12.
(Discrepancies) Suppose $\mathfrak{T}$ is an expansion subtree of $\operatorname{Tree}(R)$. Let $Q=|\text{Inner}(\mathfrak{T})|$ and $S=|\text{Leaves}(\mathfrak{T})|$ be the rod sets corresponding to those node sets. Then for $n>0$, $\operatorname{D}(n,R,S)$ is $\operatorname{C}(n,Q)$, the net number of rods in $Q$ of length $n$.
Proof.
Let $A$ be the subset of $\operatorname{Trains}(R)$ for which $m=0$ in Lemma 11, and let $B=\operatorname{Trains}(R)\setminus A$.
For trains of length $n$, we have:
$\operatorname{F}(n,R)=\operatorname{C}(n,A)+\operatorname{C}(n,B),$
where the sign of the train to its node determines the sign of the train.
Each train in $B$ starts with the train $\sigma_{1}$, with $|\sigma_{1}|\in S$, and ends with an arbitrary element of $\operatorname{Trains}(n-\text{length}(\sigma_{1}),R)$. Therefore
$\operatorname{C}(n,B)=\sum_{k\in S}(\operatorname{sign}k)\operatorname{F}(n-k,R)$
so
$\operatorname{D}(n,R,S)=\operatorname{C}(n,A).$
Finally, $\operatorname{C}(n,A)=\operatorname{C}(n,Q)$ since the trains $\tau\in A$ are precisely those that are nodes in $Q$. ∎
[…]
For Equations 3.2 and 3.3 we use the third tree in Figure 4, and note that the trains of odd length end at leaves and the trains of even length at internal nodes.
4 Equivalence
In this short section we introduce the first of the tools we need to construct our rod set algebra: canceling rod/antirod pairs. Recall that we need antirods in order to study recursions like the one in Equation (3) in the Introduction, corresponding to rod set $2,2,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}}]$. Once we allow antirods, different rod sets may have the same net train counts. For example, $\operatorname{Trains}([\phantom{a}])$ has just one element—the empty train with length 0—while $\operatorname{Trains}([1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}}])$ is the set of finite trains consisting of $1$s and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}$}s. We can pair the elements of length $n$ in $\operatorname{Trains}([1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}}])$ by switching the sign of the first rod, which guarantees that the trains in each pair have opposite signs, and shows that for any $n$ there are the same number of positive and negative trains, so $\operatorname{F}(n,[1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}}])=\operatorname{F}(n,[\phantom{a}])$ for all $n$.
The next theorem shows that this kind of equality occurs only in the obvious way.
Theorem 13.
The following assertions about rod sets $R$ and $S$ are equivalent.
  1. 1.
    $\operatorname{C}(n,R)=\operatorname{C}(n,S)$ for all $n>0$.
  2. 2.
    $\operatorname{F}(n,R)=\operatorname{F}(n,S)$ for all $n\geq 0$.
  3. 3.
    $R$ and $S$ can each be converted to the same rod set $T$ free of rod/antirod pairs by deleting some rod/antirod pairs $(k,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}})$.
  4. 4.
    $R$ can be converted to $S$ through additions and deletions of rod/antirod pairs $(k,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}})$.
Proof.
The proof is an elementary consequence of the definitions; we leave it to the reader. ∎
Definition 14.
Rod sets $R$ and $S$ are equivalent, written $R\equiv S$, when they satisfy any one (and hence all) of the conditions in the previous theorem. Each equivalence class contains a unique reduced rod set $T(R)$ that contains no rod/antirod pairs. Rod set $R$ is positive when $T(R)$ contains no antirods, is finite when $T(R)$ contains finitely many rods, and is primitive when the greatest common divisor of lengths of rods in $T(R)$ is $1$. We write $\min R$ for the length of the shortest rod in $T(R)$. When $R$ is finite we write $\|R\|$ for the number of rods in $T(R)$ and $\max R$ for the length of the longest rod in $T(R)$.
Theorem 15.
If $R\equiv R^{\prime}$ and $S\equiv S^{\prime}$ then
$R\cup S$ $\equiv R^{\prime}\cup S^{\prime}$
$R\cap S$ $\equiv R^{\prime}\cap S^{\prime}$
$RS$ $\equiv R^{\prime}S^{\prime}$
[…]
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{(RS)}}$ $\equiv{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{R}}S\equiv R{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{S}}.$ (4.1)
If $S\subseteq R$ then
$R\setminus S\equiv R\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{S}}.$
Proof.
The first six assertions follow directly from the definitions. For the last assertion, write $R$ as the disjoint union of $R^{\prime\prime}$ and $S$. Then $R\setminus S=R^{\prime\prime}$, and $R\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{S}}=R^{\prime\prime}\cup S\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{S}}\equiv R^{\prime\prime}$. ∎
Corollary 16.
The rod set $|\operatorname{Trains}(R)|$ is equivalent to the reduced rod set containing $\operatorname{F}(n,R)$ rods of length $n$ with appropriate signs.
5 Rod Set Algebra
In this section we prove that rod set expansion is well defined up to rod set equivalence and is independent of which expansion tree we use. Then we develop algebraic ways to manipulate equivalence classes of rod sets.
Figure 5 shows three expansion trees for rod sets equivalent to $[1,2]$, with each node $\tau$ labeled with the rod $|\tau|$. The three leaf sets all satisfy $|\text{Leaves}(\mathfrak{T})|\equiv[2,3,4,4,5]$ while the internal node sets satisfy $|\text{Inner}(\mathfrak{T})|\equiv[1,2,3]$. The figures suggest we can write
$1,2]\xrightarrow{\hphantom{a}[1,2,3]\hphantom{a}}[2,3,4,4,5]$
for any one of the three expansions
$\{1,2\}$ $\xrightarrow{\hphantom{a}\{1,2,21\}\hphantom{a}}\{11,12,211,212,22\},$
$\{1,2\}$ $\xrightarrow{\hphantom{a}\{1,11,111\}\hphantom{a}}\{1111,1112,112,12,2\},$
$\{1,2,3,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}}\}$ $\xrightarrow{\hphantom{a}\{1,2,3\}\hphantom{a}}\{11,12,13,1{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}},21,22,23,2{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}},31,32,33,3{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}}\}.$
Figure 5: Three trees for the expansion $1,2]\xrightarrow{\hphantom{a}[1,2,3]\hphantom{a}}[2,3,4,4,5]$.
Since $R=[1,2]$, when we construct the first two trees in Figure 5 step-by-step we have to wait to expand a node equivalent to $3$ until one appears in the tree. However, quantum mechanics suggests a way to avoid the wait: it predicts that systems (even the vacuum) may spontaneously create particle/antiparticle pairs, which usually mutually annihilate immediately. But sometimes one member of the pair decays before annihilation happens. That metaphor motivates the third expansion subtree in the figure, where we add the rod/antirod pair $3,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}}]$ to $[1,2]$ in order to expand the $3$ without having to wait for it to show up. We end up with an equivalent leaf set.
Algebraically, the metaphor corresponds to adding and subtracting $\operatorname{F}(n-3,[1,2])$ in order to write the original recursion as if it were for the rod set $1,2,3,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}}]$:
$\operatorname{F}(n,[1,2])=\operatorname{F}(n,[1,2,3,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}}])=\operatorname{F}(n-1)+\operatorname{F}(n-2)+\operatorname{F}(n-3)-\operatorname{F}(n-3),$
then using the recursion for $1,2,3,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}}]$ to replace the first three terms in the sum with four new terms each (two of these replacement terms also add to zero). After canceling equal terms what remains is the recursion for $[2,3,4,4,5]$.
If all the rods $Q$ we want to expand are already in $R$ then the rod set $|\text{Leaves}(\mathfrak{T})|$ for the expansion subtree $\mathfrak{T}$ of $\operatorname{Tree}(R)$ with $|\text{Inner}(\mathfrak{T})|=Q$ is $(R\setminus Q)\cup QR$. Theorem 15 then implies
$|\text{Leaves}(\mathfrak{T})|=(R\setminus Q)\cup QR\equiv R\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}\cup QR.$
We will soon see that up to rod set equivalence this union is the leaf rod set when expanding $R$ by $Q$, even when $Q$ is not a subset of $R$.
Definition 17.
For rod sets $R$, $Q$ and $S$ we say $R$ expands to $S$ via $Q$, and write
$R\xrightarrow{\hphantom{a}Q\hphantom{a}}S,$
if there is a rod set $R^{\prime}\equiv R$ and an expansion subtree $\mathfrak{T}$ of $\operatorname{Tree}(R^{\prime})$ for which $|\text{Inner}(\mathfrak{T})|\equiv Q$ and $|\text{Leaves}(\mathfrak{T})|\equiv S$ such that
$R^{\prime}\xrightarrow{\hphantom{a}\text{Inner}(\mathfrak{T})\hphantom{a}}\text{Leaves}(\mathfrak{T}).$
Theorem 18.
(Expansion) For rod sets $R$ and $Q$ there is a rod set $S$ such that
$R\xrightarrow{\hphantom{a}Q\hphantom{a}}S,$
and this expansion is well defined up to rod set equivalence: if $R^{\prime}\equiv R$, $Q^{\prime}\equiv Q$, and $\mathfrak{T}$ is an expansion subtree of $\operatorname{Tree}(R^{\prime})$ with $|\text{Inner}(\mathfrak{T})|=Q^{\prime}$, then $|\text{Leaves}(\mathfrak{T})|\equiv R\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}\cup QR$.
Proof.
We start by proving uniqueness. Let $\mathfrak{T}$ be any expansion subtree of $R^{\prime}\equiv R$ with internal nodes $Q^{\prime}\equiv Q$. Every node in $\mathfrak{T}$ is either the root $\Lambda$, a leaf, or an internal node in $Q^{\prime}$. Furthermore, every node in $\mathfrak{T}\setminus\{\Lambda\}$ is either a child of $\Lambda$, hence a node in $R^{\prime}$, or a child of an internal node, hence a node in $Q^{\prime}R^{\prime}$. Thus $R^{\prime}\cup Q^{\prime}R^{\prime}$ is a list of all nodes in $\mathfrak{T}\setminus\{\Lambda\}$. To list the leaves, we take out the internal nodes,
$|\text{Leaves}(\mathfrak{T})|$ $=|(R^{\prime}\cup Q^{\prime}R^{\prime})\setminus Q^{\prime}|$
$\equiv(R^{\prime}\cup Q^{\prime}R^{\prime})\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q^{\prime}}}$
$\equiv R\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}\cup QR.$
To show existence we need just one instance. Let $R^{\prime}=R\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}\cup Q\equiv R.$ Then $\mathfrak{T}=\{\Lambda\}\cup R^{\prime}\cup QR^{\prime}$ is an expansion subtree of $\operatorname{Tree}(R^{\prime})$ for which $\text{Inner}(\mathfrak{T})=Q$. ∎
Now we explore some simple consequences of this theorem that define useful operations in our rod set algebra.
Corollary 19.
(Exchange) Rod set $R$ expands via $Q$ to $S$ if and only if $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}$} expands via $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{R}$} to $S$.
Proof.
Since $QR\equiv{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}.{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{R}}$,
$R\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}\cup QR={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}\cup R\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}.{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{R}}.$
Corollary 20.
(Odd Sign Swap) Suppose $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$. Let $R^{\prime},S^{\prime}$, and $Q^{\prime}$ respectively represent $R,Q,S$ with the signs of odd length rods switched and those of even length rods kept the same. Then
$\operatorname{F}(n,R^{\prime})=(-1)^{n}\operatorname{F}(n,R)$
and
$R^{\prime}\xrightarrow{\hphantom{a}Q^{\prime}\hphantom{a}}S^{\prime}.$
Proof.
The first assertion follows from the fact that a train of odd (even) length contains an odd (even) number of rods of odd length. To prove the second, we may assume $S=R\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}\cup QR$. To show that $S^{\prime}=R^{\prime}\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q^{\prime}}}\cup Q^{\prime}R^{\prime}$ we deal separately with each of the three rod sets whose union is $S$. Reversing the odd length rods in $R$ and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}$} yields $R^{\prime}$ and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q^{\prime}}$}. The set $QR$ consists of rods of the form $|q.r|$ with $q\in Q$ and $r\in R$. This concatenation has odd length when exactly one of $q$ and $r$ has odd length, so $QR=Q^{\prime}R^{\prime}$. ∎
The next theorem collects several important equivalent ways to characterize the expansion $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$.
Theorem 21.
Let $R$, $Q$ and $S$ be rod sets. Then the following assertions are equivalent.
  1. 1.
    $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S.$
  2. 2.
    For $n>0$
    $\operatorname{D}(n,R,S)=\operatorname{C}(n,Q).$ (5.1)
  3. 3.
    Trains built from rods in $R$ parse to trains built from $S$ and trains built from $S$ with a rod from $Q$ appended, so
    $|\operatorname{Trains}(R)|\equiv|\operatorname{Trains}(S)|\cup|\operatorname{Trains}(S).Q|.$ (5.2)
  4. 4.
    For $n>0$
    $\operatorname{F}(n,R)=\operatorname{C}(n,\operatorname{Trains}(R))=\operatorname{C}(n,\operatorname{Trains}(S))+\operatorname{C}(n,\operatorname{Trains}(S).Q).$ (5.3)
Proof.
We use any convenient expansion subtree $\operatorname{Tree}(R^{\prime})$ for a rod set equivalent to $R$ and invoke Lemma 11 and Theorem 12. ∎
As a corollary we see how to compose expansions:
Corollary 22.
(Composition) Let $P,R,S$ be rod sets with $P\xrightarrow{\hphantom{a}Q_{PR}\hphantom{a}}R$,
$R\xrightarrow{\hphantom{a}Q_{RS}\hphantom{a}}S$, and $P\xrightarrow{\hphantom{a}Q_{PS}\hphantom{a}}S$. Then
$Q_{PS}\equiv Q_{PR}\cup Q_{RS}\cup Q_{PR}Q_{RS}.$
Proof.
From the theorem we have,
$|\operatorname{Trains}(P)|$ $\equiv|\operatorname{Trains}(R)\cup\operatorname{Trains}(R).Q_{PR}|$
$|\operatorname{Trains}(R)|$ $\equiv|\operatorname{Trains}(S)\cup\operatorname{Trains}(S).Q_{RS}|.$
Substituting, we obtain
$|\operatorname{Trains}(P)|$ $\equiv|\operatorname{Trains}(S)\cup\operatorname{Trains}(S).Q_{RS}|$
$\quad\quad\cup|(\operatorname{Trains}(S)\cup\operatorname{Trains}(S).Q_{RS}).Q_{PR}|.$
We distribute, view $Q_{RS}.Q_{PR}$ as one suffix, and the corollary follows. ∎
Theorem 18 shows that there is a simple formula for $S$ in terms of $R$ and $Q$. The next theorem proves that given any two of $R,Q,S$, the third is well-defined.
Theorem 23.
Any two of $R,Q,S$ uniquely determine the third so that
$R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$.
Proof.
Theorem 18 tells us how to find $S$ given $R$ and $Q$.
Given $R$ and $S$ we can use Theorem 21 to find $Q$ by computing the discrepancies $\operatorname{D}(n,R,S)$.
[…]
It’s interesting to examine edge cases. When $Q$ is empty we have
$R\xrightarrow{\hphantom{a}[\phantom{a}]\hphantom{a}}R.$
When $Q=|\operatorname{Trains}(R)|$ every node in $\operatorname{Tree}(R)$ has been expanded, there are no leaves, and
$R\xrightarrow{\hphantom{a}|\operatorname{Trains}(R)|\hphantom{a}}[\phantom{a}].$
Since $\phantom{a}].Q=[\phantom{a}]$, we have
$\phantom{a}]\xrightarrow{\hphantom{a}Q\hphantom{a}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}.$
The last example corresponds to adding rod/antirod pairs $Q/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}$ to the empty rod set and expanding everything in $Q$ by the empty rod set, leaving no leaves other than $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}$}.
In the proof of the preceding theorem we used discrepancies to find $Q$ given $R$ and $S$. In practice, we construct an expansion tree, and hence $Q$, incrementally by hand—a technique we invented well before we worked out the formal definitions of the concepts involved. Figure 6 shows how we checked the surprising result of a computer search that told us that $[2,3]$ expands to $[4,4,4,13]$ via finite $Q$.
Figure 6: Hand calculation of $2,3]\xrightarrow{\hphantom{a}[2,3,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}},5,5,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{6}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{9}},10]\hphantom{a}}[4,4,4,13].$ Internal nodes, leaves, and canceled leaves are indicated respectively by circles, squares, and slashes.
We start by expanding both the rods in $[2,3]$ since they do not appear in $S$. We draw a square around the resulting $4$ to indicate that it’s a leaf. Since $S$ needs more $4$s, and all the leaves we might expand are too big to have $4$ as a child, we add two $4/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}}$ rod/antirod pairs at the root, then expand the $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}$}s and their descendants, canceling rod/antirod pairs as we go. Note that at length $8$ there are $2$ positive rods and one antirod. We can cancel either positive rod, leading to two non-isomorphic expansion trees. We then easily read the circled internal nodes $Q=[2,3,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}},5,5,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{6}},8,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{9}},10]$ from the final picture.
Formally, the tree in the picture is not an expansion subtree of
$\operatorname{Tree}([2,3,4,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}},4,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}}])$ because many nodes are missing some children. For example, node 2 should have children 6, $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{6}$}, 6 and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{6}$} in addition to 4 and 5. Since those missing nodes occur as rod/antirod pairs that are never expanded, we safely ignore them.
Below we formalize this algorithm. Follow along in Figure 6.
Algorithm 24.
Let $R$ and $S$ be reduced rod sets with finite $S$. The following algorithm constructs a rod set $R^{\prime}\equiv R$ and a subtree $\mathfrak{T}$ of $\operatorname{Tree}(R^{\prime})$ for which $\text{Leaves}(\mathfrak{T})\equiv S$ and
$R\xrightarrow{\hphantom{a}\text{Inner}(\mathfrak{T})\hphantom{a}}S.$
The algorithm constructs $R^{\prime}$ and $\mathfrak{T}$ incrementally. At step $n$ we construct $R_{n}\equiv R$ and $\mathfrak{T}_{n}$, a subtree of $\operatorname{Tree}(R_{n})$. The leaves of $\mathfrak{T}_{n}$ are distributed among three sets:
  • $S_{n}=S\setminus(\text{Leaves}(\mathfrak{T}_{n})\cap S$), the leaves we have not yet boxed in the figure.
  • $C_{n}$, the canceled leaves, crossed out in the figure
  • $E_{n}$, the remaining leaves, eligible for expansion.
Initialize:
  • $R_{0}=E_{0}=R$,
  • $C_{0}=[\phantom{a}]$,
  • $S_{0}=S$,
  • $\mathfrak{T}_{0}$ is the expansion subtree of $\operatorname{Tree}(R_{0})$ with $\text{Inner}(\mathfrak{T}_{0})=\emptyset$,
Loop on $n$:
foo
The algorithm ends when one of the following conditions is true:
  • $S_{n}=[\phantom{a}]$. Then $R^{\prime}=R_{n}$ and $\mathfrak{T}=\mathfrak{T}_{n}$, and $Q=\text{Inner}(\mathfrak{T})$ is finite.
  • $S_{n}\subset R_{n}$. Then $R^{\prime}=R_{n}$ and $\mathfrak{T}$ is the subtree of $\operatorname{Tree}(R^{\prime})$ built by recursively expanding the nodes in $E_{n}$ and their descendants. $Q=\text{Inner}(\mathfrak{T})$ is infinite.
When neither of those conditions is met, let $s$ be one of the shortest rods in $S_{n}\setminus(S_{n}\cap R_{n})$ and $e$ one of the shortest rods in $E_{n}$. Then update as follows:
  • If $e=s$ (both length and sign), then we have found another leaf to box:
    • $E_{n+1}=E_{n}\setminus\{e\}$,
    • $S_{n+1}=S_{n}\setminus\{s\}$,
    • $R_{n+1}=R_{n}$, $C_{n+1}=C_{n}$, and $\mathfrak{T}_{n+1}=\mathfrak{T}_{n}$.
  • If $\operatorname{length}(e)<\operatorname{length}(s)$, expand $e$:
    • $\mathfrak{T}_{n+1}$ is the tree built by expanding $e$ to $e.R_{n}$ in $\mathfrak{T}_{n}$,
    • $R_{n+1}=R_{n}$, $S_{n+1}=S_{n}$,
    • $E_{n+1}$ is the reduced rod set corresponding to $(E_{n}\cup eR_{n})\setminus\{e\}$,
    • $C_{n+1}=C_{n}\cup((E_{n}\cup eR_{n}\setminus\{e\})\setminus E_{n+1})$ is the set formed by adding to $C_{n}$ the rod/antirod pairs that were canceled in reducing to form $E_{n+1}$.
  • If $\operatorname{length}(e)>\operatorname{length}(s)$ or if $e$ and $s$ have the same lengths, but different signs, add the rod/antirod pair $s/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{s}}$ to $R_{n}$, and expand $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{s}$}:
    • $\mathfrak{T}_{n+1}$ is the tree built by adding leaves $s$ and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{s}$} as children of the root in $\mathfrak{T}_{n}$ and expanding ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{s}}\rightarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{s}}.R_{n}$,
    • $R_{n+1}=R_{n}\cup\{s,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{s}}\}$,
    • $S_{n+1}=S_{n}\setminus\{s\}$,
    • $E_{n+1}$ is the reduced rod set corresponding to $E_{n}\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{s}}R_{n}$,
    • $C_{n+1}=C_{n}\cup((E_{n}\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{s}}R_{n})\setminus E_{n+1})$ is the set formed by adding to $C_{n}$ the rod/antirod pairs that were canceled in reducing to form $E_{n+1}$.
Proof.
The algorithm terminates because one of the following will happen in a finite number of steps: $\|S_{n}\|$ decreases, $\min E_{n}$ increases, or $\min S_{n}$ increases.
Even when $S$ is infinite this algorithm generates the (possibly infinite) $Q$: every $s\in S$ becomes a leaf of $\mathfrak{T}_{n}$ for some $n$, since the length of the smallest eligible rod increases until it’s greater than the length of the smallest element left in $S_{n}$. ∎
We have been careful to ensure that all our algebra works with infinite rod sets. The next theorems show some of what we can prove when we impose finiteness conditions.
Theorem 25.
If $Q$ is finite and $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$ then for $n>\max Q$, $\operatorname{Trains}(R)$ satisfies the recursion determined by $S$. Conversely, if for some $N$, $\operatorname{Trains}(R)$ satisfies the recursion determined by $S$ for $n>N$, then $Q$ is finite and $\max Q<N$.
If $R$ is also finite, then $S$ is finite and
$\max S=\max R+\max Q.$ (5.4)
Proof.
Equation 5.1 tells us that for $n>\max Q$, the discrepancy $\operatorname{D}(n,R,S)=0$.
If $R$ is finite then the longest rod in $R\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}\cup QR$ comes from $QR$. ∎
For finite $R$ and $S$ the $Q$ in $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$ need not be finite. For example, when we use Algorithm 24 to find the $Q$ that mediates the expansion of the Fibonacci rod set $[1,2]$ to the Padovan rod set $[2,3]$ we expand the $1$, add a $3/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}}$ pair, then expand the $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}$} forever. The result is
$Q=[1,2,3,4^{2},5^{3},6^{5},7^{8},8^{13},\ldots].$
The algorithm explains why the Fibonacci multiplicities appear in $Q$: expanding $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}$} forever means appending $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}$} to every internal node in $\operatorname{Tree}([1,2])$. The next theorem generalizes this observation.
Theorem 26.
Suppose $R$ and $S$ are finite rod sets and $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$. Let $m$ be the longest rod in $S\setminus R$. Then for $n>m$, $\operatorname{C}(n,Q)$ satisfies the recursion determined by $R$.
Proof.
Given $R$ and $S$, one way to construct a tree for the expansion is to add rod/antirod pairs to $R$ for the rods in $S$ that are not already in $R$, then expand everything that is not in $S$ forever. Formally, let $R^{\prime}=R\cup(S\setminus R)\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{(S\setminus R)}}\equiv R$. Then $S\subseteq R^{\prime}$, so $R^{\prime}$ is a disjoint union $S\cup U$, with $\max U=m$. Let $\mathfrak{T}$ be the expansion subtree of $\operatorname{Tree}(R^{\prime})$ for which $\text{Inner}(\mathfrak{T})$ is $U$ and all its descendants. Then $S=\text{Leaves}(\mathfrak{T})$ and $Q=\text{Inner}(\mathfrak{T})$.
$\text{Inner}(\mathfrak{T})$ is the disjoint union of the subtrees $\mathfrak{T}_{k}$ descending from the nodes $k\in U$. For each such $k$, the net number of nodes in $\mathfrak{T}_{k}$ of length $n$ is $\operatorname{F}(n-k,R)$. For each $k$, this sequence satisfies the recursion for $R$ when $n>m\geq k$, so the same is true for their sum
$\operatorname{C}(n,Q)=\sum_{k\in U}\operatorname{F}(n-k,R).$
Corollary 27.
If $R$ and $S$ are finite rod sets with $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$, then $Q$ is finite if and only if starting for some $N\geq\max S-\max R$ there are $\max R$ consecutive rod lengths $k$ for which $\operatorname{C}(k,Q)=0$.
Proof.
The forward direction is obvious. The converse follows from the fact that $\operatorname{C}(n,Q)$ satisfies the recursion determined by $R$ for $n>\max S$ so all terms after $N$ are also zero and $Q$ is finite. ∎
We close this section by using rod set algebra to study sequences that satisfy a finite recursion but don’t start with our default initial conditions. The well-known Lucas number sequence
$\mathscr{L}=2,1,3,4,7,11,18,\ldots$ (5.5)
is defined by its first two terms and the requirement that thereafter it satisfies the Fibonacci recursion determined by the rod set $[1,2]$:
$\mathscr{L}(n)=\mathscr{L}(n-1)+\mathscr{L}(n-2).$ (5.6)
If we prepend $\mathscr{L}(0)=1$ to the sequence in Equation 5.5 to form
$1,\mathscr{L}=1,2,1,3,4,7,11,18,\ldots,$ (5.7)
then Theorem 7 implies there is a rod set $R$ whose net train counts are $1,\mathscr{L}$. But that rod set is uninformative: it does not clearly connect the Lucas number sequence to the Fibonacci recursion determined by the rod set $[1,2]$. However, Theorem 21 says that the rod set $Q$ for which
$R\xrightarrow{\hphantom{a}Q\hphantom{a}}[1,2]$
is precisely the set of rods that specify when $1,\mathscr{L}$ differs from the Fibonacci sequence determined by $[1,2]$.
Therefore
[…]
and $Q=[1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}]$. Thus, also by Theorem 21, the Lucas numbers satisfy the recursion
$\mathscr{L}(n)=\operatorname{F}(n,[1,2])+\operatorname{F}(n-1,[1,2])-2\operatorname{F}(n-2,[1,2]),$ (5.8)
which expresses each Lucas number in terms of Fibonacci numbers with default initial conditions.
Theorem 21 provides more. Equation 5.3 shows a nice way to view the trains counted by $\mathscr{L}$:
$|\operatorname{Trains}(R)|\equiv|\operatorname{Trains}([1,2])|\cup|\operatorname{Trains}([1,2]).[1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}]|.$
Table 2 illustrates the Lucas trains for the first few values of $n$ as either Fibonacci trains or Fibonacci trains with an extra suffix from $Q$ .
Equation 5.3 constructs the extra trains using the rods in $Q$ as suffixes. Benjamin and Quinn[2] handle initial conditions a little differently. They put the suffix first and call it a phase. Their trains always start with a phase, so the trains in our model that don’t have a suffix would correspond to using rods from $R$ as phases as well, and some of these rods might cancel with rods from $Q$.
Length
Trains([1,2])([1,2])
[…]
Net Total
1 1 $\Lambda.1$ 2
[…]
3 111,12,21 11.1, 2.1, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}. 1$$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}$} 3
[…]
Table 2: Rod trains for the Lucas numbers
The argument in this Lucas number example relies only on Theorem 21, not on the finiteness of $S=[1,2]$ nor on the fact that the discrepancies occur only for values $n\leq\max S$. But when these finiteness conditions are met the result shows how to use our rod model to define “traditional” initial conditions.
If $S$ is a finite rod set and
$\mathscr{L}=\ell_{1},\ell_{2},\ldots$
is a sequence starting at index $1$ for which
$\ell_{n}=\sum_{k\in R}(\operatorname{sign}k)\ell_{n-k}\quad\text{for }n>\max S.$
then $\ell_{1},\cdots,\ell_{\max S}$ are the traditional initial conditions that together with $S$ determine $\mathscr{L}$.
6 Generating functions
In this section we see how our rod set model for counting trains offers a concrete combinatorial way to interpret the power series algebra that appears when we use generating functions to study recursions. In the application sections that follow we will sometimes complement combinatorial rod set algebra arguments with references to polynomials and power series.
Our algebraic operations on rod sets define interactions between two different formal power series: the rod count generating function and the recursion generating function.
Definition 28.
The rod count generating function for rod set $R$ is the formal power series
$\operatorname{\mathcal{C}}(x,R)=\sum_{k\in R}(\operatorname{sign}k)x^{k}=\sum_{n>0}\operatorname{C}(n,R)x^{n}.$
Since there are no rods with length 0, the constant term in this power series is 0. When $R$ is finite, $\operatorname{\mathcal{C}}(x,R)$ is a polynomial.
The recursion generating function for the recursion defined by $R$ is the formal power series
$\operatorname{\mathcal{F}}(x,R)$ $=\sum_{n=0}^{\infty}\operatorname{F}(n,R)x^{n}$
$=1+\operatorname{F}(1,R)x+\operatorname{F}(2,R)x^{2}+\operatorname{F}(3,R)x^{3}+\cdots\ .$
For example,
$\operatorname{\mathcal{C}}(x,[1,2])=x+x^{2},$
and
$\operatorname{\mathcal{F}}(x,[1,2])=1+x+2x^{2}+3x^{3}+5x^{4}+8x^{5}+\cdots.$
It follows directly from the definitions that
$\operatorname{\mathcal{C}}(x,R)=-\operatorname{\mathcal{C}}(x,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{R}}),$
and that
$\operatorname{\mathcal{F}}(x,R)=1+\operatorname{\mathcal{C}}(x,|\operatorname{Trains}(R)|).$
If $R\equiv R^{\prime}$ then
$\operatorname{\mathcal{C}}(x,R)=\operatorname{\mathcal{C}}(x,R^{\prime})\text{ and }\operatorname{\mathcal{F}}(x,R)=\operatorname{\mathcal{F}}(x,R^{\prime}).$
The next lemma establishes basic properties of the rod count generating function.
Lemma 29.
For rod sets $Q,R$
$\operatorname{\mathcal{C}}(x,Q\cup R)=\operatorname{\mathcal{C}}(x,Q)+\operatorname{\mathcal{C}}(x,R)$ (6.1)
and
$\operatorname{\mathcal{C}}(x,QR)=\operatorname{\mathcal{C}}(x,Q)\operatorname{\mathcal{C}}(x,R).$ (6.2)
Proof.
The first assertion is obvious. To prove the second, note that the rod set $QR$ consists of the rods $|qr|$ with $q\in Q$ and $r\in R$, which correspond to the terms in the sum that defines $\operatorname{\mathcal{C}}(x,QR)$. ∎
The algebra for rod count and recursion generating functions works smoothly because we defined the rod count generating function $\operatorname{\mathcal{C}}(n,R)$ only for $n>0$. We add or subtract this quantity from $1$ as needed, as illustrated in the next theorem.
Theorem 30.
Rod set $R$ expands to $S$ via $Q$ if and only if
$1-\operatorname{\mathcal{C}}(x,S)=(1-\operatorname{\mathcal{C}}(x,R))(1+\operatorname{\mathcal{C}}(x,Q)).$ (6.3)
Proof.
If $R$ expands to $S$ via $Q$ then Theorem 18 implies $S\equiv R\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}\cup QR$ so
$1-\operatorname{\mathcal{C}}(x,S)$ $=1-(\operatorname{\mathcal{C}}(x,R)-\operatorname{\mathcal{C}}(x,Q)+\operatorname{\mathcal{C}}(x,QR))$
$=1-(\operatorname{\mathcal{C}}(x,R)-\operatorname{\mathcal{C}}(x,Q)+\operatorname{\mathcal{C}}(x,Q)\operatorname{\mathcal{C}}(x,R))$
$=(1-\operatorname{\mathcal{C}}(x,R))(1+\operatorname{\mathcal{C}}(x,Q)).$
Conversely, this equation implies $S\equiv R\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}\cup QR$, so $R$ expands to $S$ via $Q$. ∎
Equation 6.3 shows how useful it is to define rod generating functions without specifying $\operatorname{\mathcal{C}}(0,R)$: we need both of $1\pm\operatorname{\mathcal{C}}(x,R)$. The commutativity of polynomial multiplication does not smoothly translate to the rod context, where $R$ and $Q$ play different roles. For example, the polynomial equation
$(1-x-x^{3}-x^{4})=(1-x-x^{2})(1+x^{2}),$
corresponds both to the Fibonacci expansion $1,2]\xrightarrow{\hphantom{a}[2]\hphantom{a}}[1,3,4]$ and to its exchange, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}]\xrightarrow{\hphantom{a}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}]\hphantom{a}}[1,3,4]$. For the first expansion, we interpret the right side of the equation as $\operatorname{\mathcal{C}}(x,R)=x+x^{2}$ and $\operatorname{\mathcal{C}}(x,Q)=x^{2}$, for the second as $\operatorname{\mathcal{C}}(x,R)=-x^{2}$ and $\operatorname{\mathcal{C}}(x,Q)=-x^{2}-x$.
The infinite Fibonacci expansion $1,2]\xrightarrow{\hphantom{a}[\text{evens}]\hphantom{a}}[\text{odds}]$ corresponds to the power series identity
$(1-x-x^{3}-x^{5}-\cdots)=(1-x-x^{2})(1+x^{2}+x^{4}+\cdots).$
The Exchange Corollary 19 is an immediate consequence of Theorem 30. The Odd Sign Swap Corollary 20 follows when we replace $x$ with x-x in Equation 6.3.
This theorem shows that the Expansion Theorem 18 is essentially equivalent to the standard algorithm for multiplying formal power series. The net count of the rods of length $n$ in $R\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}\cup QR$ encodes the calculation of the coefficient of $x^{n}$ in the product power series. Then the Composition Corollary 22 is just the distributive property. Algorithm 24 illustrated in Figure 6 calculates the quotient of two power series.
While working on this paper we often found that we made fewer arithmetic errors when we calculated expansions using trees rather than by multiplying and dividing polynomials or power series using the standard algorithms. However, we do not recommend teaching these methods in schools.
Corollary 31.
If $R$ is a finite rod set then the rod count generating functions for the rod sets $S$ that are expansions of $R$ by a finite $Q$ are precisely the polynomials with constant term $1$ in the ideal generated by $1-\operatorname{\mathcal{C}}(x,R)$ in the polynomial ring $\mathbb{Z}[x]$.
Proof.
Every polynomial with constant term $1$ is $1-\operatorname{\mathcal{C}}(x,R)$ for a finite rod set $R$, and conversely. If polynomial $s(x)$ has constant term $1$ and satisfies
$s(x)=q(x)(1-\operatorname{\mathcal{C}}(x,R))$
then $q(x)$ has constant term $1$ and $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$ for the rod sets $S$ and $Q$ corresponding to the polynomials $s(x)$ and $q(x)$. ∎
The next corollary is a reprise of Theorem 23.
Corollary 32.
Any two of $R,Q,S$ uniquely determine the third so that $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$.
Proof.
Using Equation 6.3 we can solve for any one of the rod count generating functions in terms of the other two. ∎
Theorem 33.
For rod set $R$
$\operatorname{\mathcal{F}}(x,R)=\frac{1}{1-\operatorname{\mathcal{C}}(x,R)},$ (6.4)
and $R$ expands to $S$ via $Q$ if and only if
$\operatorname{\mathcal{F}}(x,R)=(1+\operatorname{\mathcal{C}}(x,Q))\operatorname{\mathcal{F}}(x,S).$ (6.5)
Proof.
Since $R$ expands to $S=[\phantom{a}]$ via $Q=|\operatorname{Trains}(R)|$, and $\operatorname{\mathcal{C}}(x,[\phantom{a}])=0$, Equation 6.3 implies
$1$ $=1-\operatorname{\mathcal{C}}(x,[\phantom{a}])$
$=(1-\operatorname{\mathcal{C}}(x,R))(1+\operatorname{\mathcal{C}}(x,|\operatorname{Trains}(R)|))$
$=(1-\operatorname{\mathcal{C}}(x,R))\operatorname{\mathcal{F}}(x,R).$
The second assertion follows from Theorem 30. ∎
For example, the recursion generating function for the Fibonaccis is
$\operatorname{\mathcal{F}}(x,[1,2])=\dfrac{1}{1-x-x^{2}}.$
For the expansion $1,2]\xrightarrow{\hphantom{a}[2]\hphantom{a}}[1,3,4]$, we have
$\operatorname{\mathcal{F}}(x,[1,2])=\dfrac{1}{1-x-x^{2}}=\dfrac{1+x^{2}}{1-x-x^{3}-x^{4}}=(1+x^{2})\operatorname{\mathcal{F}}(x,[1,3,4]).$
The formal expansion of the geometric series in Equation 6.4 is
$\frac{1}{1-\operatorname{\mathcal{C}}(x,R)}$ $=1+\operatorname{\mathcal{C}}(x,R)+\operatorname{\mathcal{C}}(x,R)^{2}+\cdots$
$=1+\operatorname{\mathcal{C}}(x,R)+\operatorname{\mathcal{C}}(x,R.R)+\cdots.$
which nicely mirrors the decomposition of $\operatorname{Trains}(R)$ in Equation 1.1.
The next corollary is a reprise of Theorem 26.
Corollary 34.
Suppose $R$ and $S$ are finite rod sets and $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$. Then for $n>\max S$, $\operatorname{C}(n,Q)$ satisfies the recursion determined by $R$.
Proof.
Equations 6.3 and 6.4 imply that
$1+\operatorname{\mathcal{C}}(x,Q)=(1-\operatorname{\mathcal{C}}(x,S))\operatorname{\mathcal{F}}(x,R).$
The right side of the equation is a finite sum of shifts of the net train count generating function for $R$. For $n>\max S$, there are no more new shifts, and the coefficients of $\operatorname{C}(n,Q)$ satisfy the recursion for $R$. ∎
The following theorem is a reprise of Theorem 7.
Theorem 35.
Every formal power series with constant term $1$ is the recursion generating function $\operatorname{\mathcal{F}}(x,R)$ for a rod set $R$ unique up to rod/antirod equivalence.
Proof.
Let $\operatorname{F}(x)=1+\sum_{k=1}^{\infty}a_{k}x^{k}$. We calculate the formal power series inverse
$\frac{1}{\operatorname{F}(x)}=1+\sum_{k=1}^{\infty}b_{k}x^{k}$
Then we let $R$ be the rod set with $b_{k}$ rods (or antirods) of length $k$ . ∎
We end this section with a well-known theorem on rational generating functions[7]:
Theorem 36.
The net train counts for a rod set $R$ eventually satisfy the recursion for a finite rod set $S$ if and only if $\operatorname{\mathcal{F}}(x,R)$ is a quotient of polynomials.
Proof.
Suppose $\operatorname{C}(n,R)$ eventually satisfies the recursion determined by a finite rod set $S$. Let $Q$ be the rod set for which
$R\xrightarrow{\hphantom{a}Q\hphantom{a}}S.$
Since $Q$ specifies the discrepancies in this expansion, Theorem 25 implies it is finite, so
$\operatorname{\mathcal{F}}(x,R)=\frac{1+\operatorname{\mathcal{C}}(x,Q)}{1-\operatorname{\mathcal{C}}(x,S)}$ (6.6)
is a quotient of polynomials.
Conversely, if the numerator and denominator in Equation 6.6 are polynomials, then $Q$ and $S$ are finite and the net train counts for $R$ satisfy the recursion for $S$ when $n>\max Q$. ∎
The generating function for the Lucas number sequence 5.7 with a prepended constant term $1$ is
$\frac{1+x-2x^{2}}{1-x-x^{2}}=1+2x+x^{2}+3x^{3}+4x^{4}+7x^{5}+\cdots.$
To find the generating function for the Lucas number sequence with traditional initial conditions 5.5, we can subtract $1$ and divide by $x$:
$\frac{2-x}{1-x-x^{2}}=2+x+3x^{2}+4x^{3}+7x^{4}+\cdots.$
7 Duality and its Applications
In this section we build one more operation for our rod set algebra. When we know how to expand $R$ to $S$, duality tells us how to expand $S$ to $R$. Duality is especially useful for exploring relationships between combinatorial identities and compositions, and we finish by deriving some Fibonacci and Padovan examples.
Definition 37.
The dual of rod set $Q$ is
$Q^{*}=|\operatorname{Trains}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}})|.$
Theorem 38.
If
$R\xrightarrow{\hphantom{a}Q\hphantom{a}}S,$
then
$S\xrightarrow{\hphantom{a}Q^{*}\hphantom{a}}R.$
Proof.
Figure 7 shows an expansion of $\operatorname{Tree}(R\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}\cup QR)$. Each node in the picture is actually a set of siblings corresponding to the rods in the rod set label. The powers of $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}$} are the internal nodes.
Figure 7: Expanding $\operatorname{Tree}(R\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}\cup QR)$ via $\operatorname{Trains}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}$
The union of the internal nodes is $\operatorname{Trains}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}})$. The leaves on the left are ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}^{k}R$ starting at $k=0$, while the leaves on the right are ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}^{k-1}QR$, starting at $k=1$. Equation 15 shows that all but the leaves in $R$ cancel. ∎
Corollary 39.
For rod set $Q$,
$Q^{**}\equiv Q.$
Proof.
If $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$, then $S\xrightarrow{\hphantom{a}Q^{*}\hphantom{a}}R$ so $R\xrightarrow{\hphantom{a}Q^{**}\hphantom{a}}S$. Then Theorem 18 implies $Q^{**}\equiv Q$. ∎
It is easy to write the rod generating function for the dual $Q^{*}$ in terms of the rod generating function for $Q$.
Theorem 40.
For rod set $Q$, we have
$1+\operatorname{\mathcal{C}}(x,Q^{*})=\frac{1}{1+\operatorname{\mathcal{C}}(x,Q)}.$
Proof.
The dual $Q^{*}$ counts the trains of $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}$} so
$1+\operatorname{\mathcal{C}}(x,Q^{*})=\operatorname{\mathcal{F}}(x,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}})=\frac{1}{1-\operatorname{\mathcal{C}}(x,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}})}=\frac{1}{1+\operatorname{\mathcal{C}}(x,Q)}.$
When $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$, Theorem 18 shows how to compute $S$ given $R$ and $Q$, and Theorem 23 tells us that any two of $R,Q,S$ determines the third. With duality, it is straightforward, albeit a bit messy, to find explicit rod set formulas for $Q$ in terms of $R$ and $S$ and for $R$ in terms of $Q$ and $S$.
Duality is particularly useful when expanding by a rod set $Q$ with just a single rod $q$. Then $\operatorname{Trains}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}})$ contains rods with lengths all the multiples of the length of $q$. Then duality reveals an interesting connection between identities involving sums and those involving compositions.
For example, recall that we established the Fibonacci identities in Equations 3.2 and 3.3 using the third expansion tree in Figure 4 together with Theorem 21. We started with the infinite expansion
$1,2]\xrightarrow{\hphantom{a}|\operatorname{Trains}(2)|\hphantom{a}}[1,3,5,7,\ldots],$ (7.1)
and deduced the identities
$1+2+5+13+34+\ \cdots\ +\operatorname{F}(2n)$ $=\operatorname{F}(2n+1),$
$1+3+8+21+\cdots+\operatorname{F}(2n-1)$ $=\operatorname{F}(2n)-1.$
Now we use Theorem 38 to dualize Equation 7.1:
$1,3,5,7,\ldots]\xrightarrow{\hphantom{a}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}]\hphantom{a}}[1,2].$ (7.2)
Theorem 12 implies the train counts for the rod set $1,3,5,7,\ldots]$ satisfy the Fibonacci recursion when $n>2$. The train counts for the rod set of odd numbers are the number of compositions, $\mathscr{O}(n)$, of $n$ into odd parts. Since $\mathscr{O}(1)=\mathscr{O}(2)=1$, the sequence $\mathscr{O}(n)$ is the Fibonacci sequence with these initial conditions, hence $\mathscr{O}(n)=\operatorname{F}(n-1,[1,2])$. For example, there are $\operatorname{F}(4,[1,2])=5$ rod trains of length $5$ built from odd length rods: $5,311,131,113,11111$.
We can create similar relationships between identities and compositions by starting with a rod set $R$ containing $q$ and expanding along all the edges in $\operatorname{Tree}(R)$ with label $q$. That expansion has internal nodes $Q=|\operatorname{Trains}([q])|$. The leaves attached to these internal nodes form a union of arithmetic progressions with difference $q$. The dual of the expansion reveals an identity about compositions.
Here is a Padovan example, expanding $[2,3]$ by $|\operatorname{Trains}(2)|$:
$2,3]\xrightarrow{\hphantom{a}|\operatorname{Trains}(2)|\hphantom{a}}[3,5,7,9,\ldots],$
with dual
$3,5,7,9,\ldots]\xrightarrow{\hphantom{a}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}]\hphantom{a}}[2,3].$
Letting $P(n)=\operatorname{F}(n,[2,3])$, the original expansion leads to the identities
$\begin{array}[]{ll}P(0)+P(2)+P(4)+\cdots+P(2n)&=P(2n+3)\\ P(1)+P(3)+P(5)+\cdots+P(2n+1)&=P(2n+4)-1.\end{array}$
From the dual we see that the number of compositions of $n\geq 3$ into odd parts greater than or equal to $3$ satisfies the Padovan recursion. Examining the initial conditions tells us that the number of such compositions is P(n3)P(n-3).
If instead we expand the Padovans by $|\operatorname{Trains}([3])|$, we discover
$\begin{array}[]{lll}P(0)+P(3)+P(6)+\cdots+P(3n)&=P(3n+2)\\ P(1)+P(4)+P(7)+\cdots+P(3n+1)&=P(3n+3)-1\\ P(2)+P(5)+P(8)+\cdots+P(3n+2)&=P(3n+4).\end{array}$
Then the dual shows that the number of compositions of $n\geq 4$ into parts congruent to $2$ mod $3$ satisfies the Padovan recursion. Examining the initial conditions tells us that the number of such compositions is P(n2)P(n-2). The reader is invited to find more such identities.
[…]
The expansion
$1,2,3,\ldots]\xrightarrow{\hphantom{a}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}}]\hphantom{a}}[1,1]$ (7.3)
connects two well-known theorems. The forward expansion tells us that the total number of compositions of $n$ is $2^{n-1}$, while the dual gives the identity
$1+2+2^{2}+\cdots+2^{n-1}=2^{n}-1.$ (7.4)
We end this section with a more complicated application using an expansion of two copies of an arithmetic sequence,
$S=[2,2,5,5,8,8,11,11,\ldots]\xrightarrow{\hphantom{a}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}}]\hphantom{a}}[2,2,3],$
with dual
$R=[2,2,3]\xrightarrow{\hphantom{a}|\operatorname{Trains}([3])|\hphantom{a}}[2,2,5,5,8,8,11,11,\ldots].$
Since $R=[2,2,3]$ is an expansion of the Fibonacci rod set $[1,2]$ by the singleton rod set $[1]$, Corollary 22 implies $1,2]\xrightarrow{\hphantom{a}\hphantom{a}}S$ has discrepancies $3,6,9,\ldots]\cup[1,4,7,\ldots]$. The discrepancies are the rods with lengths $n\equiv 0,1\pmod{3}$, which leads to the following identities for the Fibonacci sequence $\operatorname{F}(n)=\operatorname{F}(n,[1,2])$:
$\begin{array}[]{ll}\operatorname{F}(0)+\operatorname{F}(3)+\operatorname{F}(6)+\cdots+\operatorname{F}(3n)&=\dfrac{\operatorname{F}(3n+2)}{2},\\[8.0pt] \operatorname{F}(1)+\operatorname{F}(4)+\operatorname{F}(7)+\cdots+\operatorname{F}(3n-2)&=\dfrac{\operatorname{F}(3n)-1}{2},\\[8.0pt] \operatorname{F}(2)+\operatorname{F}(5)+\operatorname{F}(8)+\cdots+\operatorname{F}(3n-1)&=\dfrac{\operatorname{F}(3n+1)-1}{2}.\end{array}$
So we can also use rod set algebra to derive identities that require fractions!
8 Rod Sets and Binomial Coefficients
It’s well known that the Fibonacci numbers are sums of diagonals in Pascal’s triangle222a.k.a. Yang Hui’s, Tartaglia’s, et.al’s triangle. More generally, the net train counts $\operatorname{F}(n,R)$ also appear as sums of diagonals when $R$ consists of two positive rods, which leads to some nice binomial identities. Even more generally, when $R$ is positive and finite we can (but won’t) systematically compute $\operatorname{F}(n,R)$ as a sum of multinomial coefficients on hyperplanes in high dimensional analogues of the triangle.
Definition 41.
Let $R$ be a rod set and $\tau\in\operatorname{Trains}(R)$. The underlying partition of $\tau$ is the multiset of rods in $\tau$.
For example, if $R=[3,4]$ and $\tau=34443$ then the underlying partition of $\tau$ is $\{3,3,4,4,4\}$.
Theorem 42.
Let $R=[a,b]$ be a positive rod set. Then
$\operatorname{F}(n,R)=\sum_{\begin{subarray}{c}ia+jb=n\\ i,j\geq 0\end{subarray}}\binom{i+j}{j}$. (8.1)
Proof.
For each pair of integers $i,j\geq 0$ with $ia+ja=n$, let $U$ be the multiset containing $i$ rods of length $a$ and $j$ rods of length $b$. A length $n$ train in $\operatorname{Trains}(R)$ with underlying partition $U$ consists of $i+j$ rods for some pair $i,j$. We count the number of these trains by starting with $i+j$ places and then choosing in $\binom{i+j}{j}$ ways the places for a rod of length $b$. We sum over the possible partitions to calculate $\operatorname{F}(n,R)$. ∎
If $R$ contains more than two rods, we can use the same logic to write $\operatorname{F}(n,R)$ as a sum of multinomial coefficients; the linear constraint will just be an equation with more variables. When $R$ contains antirods we can adjust the signs in the summation.
If $R=[a,b]$ is primitive and positive, then the Chinese Remainder Theorem implies that integer solutions $(i,j)$ to $ia+jb=n$, the condition on the sum in Equation 8.1, are congruent mod $a\times b$. We care only about those for non-negative $i$ and $j$.
To visualize the Pascal triangle diagonals, we order the terms in the sum by increasing $j$, starting with the fewest possible rods of length $b$. At each step we replace $b$ rods of length $a$ with $a$ rods of length $b$ to build the next partition, ending with the partition with the most rods of length $b$.
For example, if $R=[3,5]$ and $n=70$, we start with the underlying partition $\{3^{20},5^{2}\}$ and then move to the underlying partition $\{3^{15},5^{5}\}$, and continue until we have the sum
$\operatorname{F}(70,[3,5])=\binom{22}{2}+\binom{20}{5}+\binom{18}{8}+\binom{16}{11}+\binom{14}{14}=63862.$ (8.2)
As we move from term to term in the sum, the number of rods in the underlying partition decreases by ba=2b-a=2, and the number of rods of length $b$ increases by $a=3$. The constant increases and decreases lead to the terms in the sum appearing on parallel diagonals in the triangle. Figure 8 shows the diagonals for the Fibonacci and Padovan rod sets. When $R=[a,b]$ is primitive, the first diagonal with two terms is
$\operatorname{F}(a\times b,[a,b])=\binom{a}{0}+\binom{b}{b}=\binom{a}{a}+\binom{b}{0}=2.$
The two ways of writing this sum correspond to the dark diagonals in Figure 8. The first sum counts rods of length $b$ and the second rods of length $a$. Above the dark diagonal, the diagonal for $\operatorname{F}(n,[a,b])=0$ is empty (dashed in the figure) unless $n$ is a multiple of $a$ or $b$. Below the dark diagonal the train counts are positive. In practice, it’s hard to draw the diagonals for large $a$. Hopkins [8, Appendix A] shows similar diagrams.
Figure 8: Train counts for the Fibonaccis, $[1,2]$, and Padovans, $[2,3]$, as diagonal sums in Pascal’s triangle
Theorem 42 leads to some nice binomial identities. The train counts for $R=[1,1]$ are powers of two. The corresponding “diagonals” are the rows of the triangle, providing another proof of the familiar identity
$\operatorname{F}(n,[1,1])=\sum_{i=0}^{n}\binom{n}{i}=2^{n}.$
When we allow antirods in $[a,b]$ and adjust the signs of the binomial coefficients appropriately we see more nice identities.
For $R=[1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}}]$ the net train counts are all zero for $n\geq 1$, hence
$\operatorname{F}(n,[1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}}])=\sum_{i=0}^{n}(-1)^{i}\binom{n}{i}=0,$
where $i$ represents the number of antirods in the underlying partition.
When $R=[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}]$ the total train counts are the Fibonacci numbers, but the net train counts have period $3$. Checking the signs carefully leads to the identity
[…]
Here $i$ is the number of rods of length $2$ in the underlying partition and nin-i is the total number of rods. Since both rods are antirods, the sign depends on the parity of the number of rods.
9 Periodicity
At the end of Section 7 we looked at expansions where $\|Q\|=1$; here we’ll explore expansions where $\|S\|=1$. These seemingly similar conditions are conceptually different: the latter leads to periodicity. The periodicity in Equation 8.3 corresponds to the expansion
$1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}]\xrightarrow{\hphantom{a}[1]\hphantom{a}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}}].$ (9.1)
The net train counts for the rod set $1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}]$ start 1,1,0,1,1,01,1,0,-1,-1,0 and then repeat with period $6$, as shown in Table 1 and Equation 3.1. The periodicity is equivalent to the expansion
$1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}]\xrightarrow{\hphantom{a}[1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}}]\hphantom{a}}[6].$ (9.2)
Theorem 25 says that that when $n>4$, the net train counts $\operatorname{F}(n,[1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}])$ satisfy the recursion $\operatorname{F}(n)=\operatorname{F}(n-6)$ determined by rod set $[6]$.
This example generalizes easily.
Definition 43.
The net train counts for rod set $R$ are eventually periodic with period $p>0$ when there is an $N\geq 0$ such that
$\operatorname{F}(n+p,R)=\operatorname{F}(n,R)\text{ for all }n\geq N.$
When $N=0$ we say the sequence is periodic.
Any positive multiple of a period is a period. Sometimes we will use the period for the least period.
The following theorem characterizes rod sets with eventually periodic net train counts.
Theorem 44.
(Periodicity) The net train counts for rod set $R$ are eventually periodic with period $p$ if and only if $R$ expands to $[p]$ via a finite $Q$. The net train counts $\operatorname{F}(n,R)$ are periodic for $n>\max Q$.
Proof.
This is the special case of Theorem 25 when $S$ contains just a single positive rod. ∎
Expansions to a single antirod are related to periodicity. For example, expanding the $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}$} in Equation 9.1 leads to the familiar period $6$ expansion,
$1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}]\xrightarrow{\hphantom{a}[1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}}]\hphantom{a}}[6].$
We generalize in the following theorem.
Theorem 45.
If $R$ expands to $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}$}] via $Q$, then $R$ expands to $[2k]$ via $Q\cup[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}}]\cup Q{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}}$, and $\operatorname{F}(n,R)$ is eventually periodic with period $2k$.
Proof.
Since $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}$}] expands via $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}$}] to $[2k]$, Corollary 22 implies $R$ expands to $[2k]$ via $Q\cup[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}}]\cup Q{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}}$. ∎
There are many rod sets $R$ with eventually periodic net train counts, since every sequence of integers counts the net trains for some rod set $R$. But such $R$ are rarely finite. The following characterization of periodic finite rod sets includes a proof that eventually periodic finite rod sets are actually periodic.
Theorem 46.
The following assertions about finite rod set $R$ are equivalent.
  1. 1.
    The net train counts for $R$ are eventually periodic.
  2. 2.
    $\operatorname{F}(n,R)$ is bounded.
  3. 3.
    $\operatorname{F}(n,R)$ contains two identical sequences of length $\max R$.
  4. 4.
    The net train counts $\operatorname{F}(n,R)$ are periodic starting at $n=0$.
  5. 5.
    The polynomial $1-\operatorname{\mathcal{C}}(x,R)$ divides $1-x^{p}$ for some $p$.
  6. 6.
    $1-\operatorname{\mathcal{C}}(x,R)$ is a product of distinct cyclotomic polynomials.
  7. 7.
    The recursion generating function for $R$ is a finite sum of geometric series of the form
    $\sum_{n=0}^{\infty}\operatorname{F}(n,R)x^{n}$ $=\frac{q(x)}{1-x^{m}}$ (9.3)
    $=q(x)\left(1+{x^{m}}+{x^{2m}}+\cdots\right),$
    where
    $q(x)=\sum_{k=0}^{m-1}\operatorname{F}(k,R)x^{k}\ .$
Proof.
Clearly (1) implies (2) even without the finiteness hypothesis.
The pigeonhole principle shows that (2) implies (3): there are only finitely many bounded sequences of length $\max R$.
To show (3) implies (4), we start with two identical sequences of length $m=\max R$ starting at $a$ and $a+p$.
The recursion shows that $\operatorname{F}(a+m)=\operatorname{F}(a+p+m)$ and then by induction that for all $n\geq a$, $\operatorname{F}(n)=\operatorname{F}(n+p)$. Thus $\operatorname{F}(n)$ is eventually periodic with period $p$.
Now we need to show that $\operatorname{F}(n)$ is also periodic for $n<a$, which we can do by “running the recursion backwards” to compute earlier values of $n$. We start with
$\operatorname{F}(a+m-1)$ $=\sum_{k\in R}\operatorname{F}(a+m-1-k)$
$=\operatorname{F}(a-1)+\sum_{k\in R,k<m}\operatorname{F}(a+m-1-k),$
and compute
$\operatorname{F}(a-1)$ $=\operatorname{F}(a+m-1)-\sum_{k\in R,k<m}\operatorname{F}(a+m-1-k)$
$=\operatorname{F}(a+p+m-1)-\sum_{k\in R,k<m}\operatorname{F}(a+p+m-1-k)$
$=\operatorname{F}(a+p-1).$
We continue by induction to compute earlier terms and conclude that starting at $n=0$, $\operatorname{F}(n,R)$ is periodic with period $p$. Clearly (4) implies (1), so the first four assertions are equivalent.
The last three equivalences involve the rod count and recursion generating functions for $R$.
Theorem 44 says periodicity is equivalent to $R$ expanding to the one element rod set $[p]$, for which $\operatorname{\mathcal{C}}(x,[p])=x^{p}$. When $R$ is finite Theorem 30 then implies (1) and (5) are equivalent.
The equivalence of (5) and (6) follows from the fact that the complete factorization of $x^{p}-1$ into irreducible polynomials is
$x^{p}-1=\prod_{d|m}\Phi_{d}(m)$ (9.4)
where $\Phi_{d}(x)$ is the cyclotomic polynomial whose roots are the primitive $d$th roots of unity[4].
The equivalence of (7) is Theorem 33. ∎
Repeated cyclotomic factors lead to a rod set that isn’t periodic. For example, if $1-\operatorname{\mathcal{C}}(x,R)=(1-x)^{2}=1-2x+x^{2}$ then $R=[1,1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}]$. The expansion
$1]\xrightarrow{\hphantom{a}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}}]\hphantom{a}}[1,1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}]$
has dual
$1,1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}]\xrightarrow{\hphantom{a}|\operatorname{Trains}[1]|\hphantom{a}}[1].$
Here $Q=|\operatorname{Trains}[1]|$ is infinite and the net train count sequence $\operatorname{F}(n,[1,1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}])=n+1$ is unbounded.
In Section 7 on Duality, we discussed examples where the rod lengths were infinite arithmetic progressions. Some rod sets where the lengths are finite arithmetic progressions also provide examples of periodicity.
The expansion
[…]
corresponding to the factorization
$1-x^{m}=(1-x)(1+x+x^{2}+\cdots+x^{m-1}),$
tells us that
$\operatorname{F}(n,[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}},\ldots,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{m-1}}])$ $=\begin{cases}1&\text{ if }n\equiv 0\pmod{m}\\ -1&\text{ if }n\equiv 1\pmod{m}\\ 0&\text{ otherwise}.\end{cases}$
Similar results follow from the expansions
$[1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}},3,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}},\ldots,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2m}}]$ $\xrightarrow{\hphantom{a}[1]\hphantom{a}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2m+1}}]$
$\text{and }[1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}},3,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}},\ldots,2m-1]$ $\xrightarrow{\hphantom{a}[1]\hphantom{a}}[2m].$
When $m$ has multiple prime factors the polynomial $1-x^{m}$ can lead to interesting periodic rod sets. For example,
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}},3,4,5,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{7}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{8}}]\xrightarrow{\hphantom{a}\hphantom{a}}[30].$
We encourage the reader to check this result using Algorithm 24. Theorem 45 says we can stop when we’ve expanded to $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{15}$}.
In the examples of periodic rod sets that we’ve seen so far, there has been at most one rod of each length, but there are periodic rod sets that contain multiple rods of the same length. The first such rod set, $R$, comes from factoring $\Phi(105)$. It has period $105$ with $\max R=48=\varphi(105)$ and two rods each of lengths $7$ and $41$.
10 Expandability and Scaling
From now on we will often work with expansions where all rod sets are finite. To save repeatedly typing that:
Definition 47.
The expansion $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$ is a finite expansion when all three rod sets $R$, $Q$ and $S$ are finite.
In this section we study the existence of finite expansions when we specify only the lengths of the rods in $S$, not their multiplicities. After some general definitions we focus on expansions to one or two rod lengths. Our theorems depend on interesting conditions on the net train counts of $R$. They lead to combinatorial proofs about divisibility in Fibonacci and Lucas sequences, to more identities, and to a combinatorial contribution to factoring Borwein trinomials.
As usual, we will write $\operatorname{F}(n)$ for $\operatorname{F}(n,R)$ when there is no potential ambiguity.
Definition 48.
A shape is a set of rod lengths, usually written as an ordered list in angle brackets. When shape $H$ is finite we write $|H|$ for its cardinality. The shape of rod set $R$, $\operatorname{Shape}(R)$, is the set of lengths of rods in the reduced rod set $R^{\prime}$ equivalent to $R$; if $R$ is finite, so is its shape. For each length $\ell\in\operatorname{Shape}(R)$, the multiplicityof $\ell$ is $\operatorname{C}(\ell,R^{\prime})\neq 0$, the net count of rods of that length in $R^{\prime}$. The rods in $R^{\prime}$ of length $\ell$ are antirods just when the multiplicity is negative.
We can rewrite the recursion for $R$ using shape and multiplicities:
$\operatorname{F}(n,R)$ $=\ \quad\sum_{r\in R}\ (\operatorname{sign}r)F(n-r,R)$
$=\sum_{\ell\in\operatorname{Shape}(R)}\operatorname{C}(\ell,R)\operatorname{F}(n-\ell,R).$ (10.1)
For example, if $R=[1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}},2,2,2,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}},3]$, then $\operatorname{Shape}(R)=\langle 2,3\rangle$ and the multiplicities of 2 and 3 are 3 and 1-1, respectively. When a multiplicity is negative we use the absolute value of the multiplicity to count the number of copies. In this example $R\equiv[2^{3},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}}^{1}]$.
Definition 49.
Let $R$ be a rod set and $H$ a shape. Then $R$ is expandable to $H$ when there is a finite expansion of $R$ to a rod set with shape $H$. Then $H$ is necessarily finite with cardinality $n$. Then we say that $R$ is $n$-expandable.
Note that any rod set expands trivially to its shape via the empty $Q$.
Our goal is to characterize finite expansions of a rod set $R$ to shapes $H$ of cardinality 1 and 2 in terms of conditions on the net train counts $\operatorname{F}(n,R)$ for $n\leq\max H$.
For example, the expansion
[…]
corresponding to the polynomial factorization
$1+4x^{4}=(1+2x+2x^{2})(1-2x+2x^{2}),$
[…]
$\operatorname{F}(n,R)=1,-2,2,0,-4,8,-8,0,16,-32,32,0,-64,128,-128,0\ldots.$ (10.2)
Since $\max Q=2$, Theorem 25, tells us that the $\operatorname{D}(n,R,[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}}^{4}])$ must vanish for $n\geq 3$. In particular, $\operatorname{F}(3)+4F(3-4)=\operatorname{F}(3)=0$ and $\operatorname{F}(4)+4F(0)=-4+4=0$.
If $\max S-\max R=\max Q$ were greater, then we’d have more discrepancies that would need to be 0. Similar train count conditions will be the key to our theorems.
Theorem 50.
There is a finite expansion of rod set $R$ to a rod set $S$ with shape $\langle a\rangle$ if and only if $\operatorname{F}(a-i,R)=0$ for $i=1,2,\ldots,\max R-1$ and $\operatorname{F}(a,R)\neq 0$. Then $\operatorname{C}(a,S)=\operatorname{F}(a,R)$.
Proof.
Suppose $\operatorname{Shape}(S)=\langle a\rangle$ and $Q$ is finite in the expansion $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$.
Theorem 25 implies $a=\max R+\max Q$, so the discrepancies measuring the ways the counts $\operatorname{F}(n,R)$ fail to satisfy the recursion for $S$ are 0 for $n>\max Q=a-\max R$. Thus for positive $i<\max R$
$\operatorname{F}(a-i,R)$ $=\operatorname{F}(a-i,R)-\operatorname{C}(a,S)\operatorname{F}(a-i-a,R)$
$=\operatorname{D}(a-i,R,S)$
$=0.$
Now consider the discrepancy at $a$:
$\operatorname{D}(a,R,S)=\operatorname{F}(a,R)-\operatorname{C}(a,S)\operatorname{F}(a-a,R)=0.$ (10.3)
Since $\operatorname{F}(a-a,R)=1$ we have $\operatorname{C}(a,S)=\operatorname{F}(a,R)\neq 0$. When that train count is negative, $a$ is an antirod.
To prove the converse, suppose $\operatorname{F}(a-i,R)=0$ for $i=1,2,\ldots,\max R-1$. Let $Q$ be the rod set that expands $R$ to the rod set $S$ with $\operatorname{F}(a,R)$ copies of $a$, with $a$ an antirod if $\operatorname{F}(a,R)$ is negative. The assumption on the net train counts of $R$ say that the discrepancy $\operatorname{C}(n,Q)=\operatorname{D}(n,R,S)=0$ when $a-\max R<n<a$. Setting $\operatorname{C}(a,S)=\operatorname{F}(a,R)$ guarantees discrepancy 0 at $n=a$. Then Corollary 27 implies $Q$ is finite. ∎
Next we apply similar discrepancy techniques to the more complex characterization of 2-expansions. Finite expansions of a finite $R$ to $S$ with shape $\langle a,b\rangle$ correspond to recursion identities of the form
$\operatorname{F}(n,R)=\operatorname{C}(a,S)\operatorname{F}(n-a,R)+\operatorname{C}(b,S)\operatorname{F}(n-b,R)\quad\quad\text{for }n>\max Q$
and to trinomial multiples of the polynomial $1-\operatorname{\mathcal{C}}(x,R)$.
For example there is a finite expansion of the Fibonacci rod set $[1,2]$ to $8^{48},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{12}}^{7}]$, corresponding to the Fibonacci identity
$\operatorname{F}(n)=48\operatorname{F}(n-8)-7\operatorname{F}(n-12),$
for large enough $n$, and to the factorization
$1-48x^{8}+7x^{12}=q(x)(1-x-x^{2}),$
where
$q(x)=1+x+2x^{2}+3x^{3}+5x^{4}+8x^{5}+13x^{6}+21x^{7}-14x^{8}+7x^{9}-7x^{10}.$
Later we will show how to find this and all similar two-term Fibonacci identities.
In Theorem 50 we saw that 1-expansions of $R$ correspond to sequences of 0’s in $\operatorname{F}(n,R)$. For 2-expandability the correspondence is to divisibility among net train counts. For example, we will see soon that the expansion above of $[1,2]$ to $8^{48},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{12}}^{7}]$ is equivalent to the fact that
$48\operatorname{F}(3,[1,2])=48\times 3=144=\operatorname{F}(11,[1,2]).$
The Padovan expansion of $[2,3]$ to $4^{3},13]$ in Figure 6 corresponds to the adjacent pair of scalings by a factor of 3:
$3\operatorname{F}(7,[2,3])$ $=3\times 3=\ 9=\operatorname{F}(11,[2,3])$
$3\operatorname{F}(8,[2,3])$ $=3\times 4=12=\operatorname{F}(12,[2,3])$
in the Padovan sequence
$P(n)=1,0,1,1,1,2,2,\textbf{3},\textbf{4},5,7,\textbf{9},\textbf{12},16,21,28,\ldots$
To prove our general theorem on $2$-expandability, we formally define scaling.
Definition 51.
The finite rod set $R$ scales at counts $(c,b)$ when there is a unique integer $\alpha$ such that
$\operatorname{F}(b-i,R)=\alpha\operatorname{F}(c-i,R)$ (10.4)
for $1\leq i<\max R$. Then we say $R$ scales by $\alpha$ at $(c,b)$.
The requirement that $\alpha$ be unique exclude the degenerate case where all the Equations 10.4 are of the form $0=\alpha\times 0$.
The previous examples show that $[1,2]$ scales by $3$ at $(4,12)$ and $[2,3]$ scales by $3$ at $(9,13)$.
For a larger example including antirods we revisit the Padovan expansion
$2,3]\xrightarrow{\hphantom{a}Q=[2,3,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}}^{2},5^{2},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{6}},8,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{9}},10]\hphantom{a}}[4^{3},13].$
The Exchange Corollary 19 implies
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}\xrightarrow{\hphantom{a}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}}]\hphantom{a}}[4^{3},13].$ (10.5)
The net train counts for $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}$} are
$\operatorname{F}(n,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}})=1,0,-1,-1,3,0,-3,-3,9,0,-9,-9,27,1,-27,-28,80,6,-81,\ldots.$
In the following table the the third line shifts the net train counts in the second by $4$. Then the second line is 3 times the third between $n=4$ and $n=12$.
$\begin{array}[]{rrrrrrrrrrrrrrrrrr}n:&0&1&2&3&4&5&6&7&8&9&10&11&12&13&\ldots\\ \operatorname{F}(n):&1&0&-1&-1&3&0&-3&-3&9&0&-9&-9&27&1&\ldots\\ \operatorname{F}(n-4):&&&&&1&0&-1&-1&3&0&-3&-3&9&0$$\ldots\\ \end{array}$
Therefore $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}$} scales by $3$ at $(4,13)$.
The next theorem shows that scaling is the heart of $2$-expandability.
Theorem 52.
There is a finite expansion of rod set $R$ to a rod set $S$ with shape $\langle a,b\rangle$ if and only if there is a nonzero integer $\alpha$ such that $R$ scales by $\alpha$ at (ba,b)(b-a,b), and $\operatorname{F}(b,R)\neq\alpha\operatorname{F}(b-a,R)$. Then
$\operatorname{C}(a,S)$ $=\alpha,$ (10.6)
$\operatorname{C}(b,S)$ $=\operatorname{F}(b,R)-\alpha\operatorname{F}(b-a,R).$ (10.7)
Proof.
Suppose there is a finite expansion $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$ with $\operatorname{Shape}(S)=\langle a,b\rangle$. Theorem 25 implies $b=\max R+\max Q$, so the discrepancies measuring the ways the counts $\operatorname{F}(n,R)$ fail to satisfy the recursion for $S$ are 0 for $n>b-\max R$. In the range $b-\max R<n<b$, we have $\operatorname{F}(n-b,R)=0$.
At n=bin=b-i for positive $i<\max R$ we compute
$\operatorname{D}(b-i,R,S)$ $=\operatorname{F}(b-i,R)-\operatorname{C}(a,S)\operatorname{F}(b-a-i)-\operatorname{C}(b,S)\operatorname{F}(b-b-i)$
$=\operatorname{F}(b-i,R)-\operatorname{C}(a,S)\operatorname{F}(b-a-i)$
$=0.$
Thus for $i=1,2,\ldots,\max R-1$
$\operatorname{F}(b-i,R)=\operatorname{C}(a,S)\operatorname{F}(b-a-i,R),$ (10.8)
so $R$ scales by $\alpha=\operatorname{C}(a,S)$ at (ba,b)(b-a,b).
Now consider the discrepancy at $b$:
$\operatorname{D}(b,R,S)=\operatorname{F}(b,R)-\operatorname{C}(a,S)\operatorname{F}(b-a)-\operatorname{C}(b,S)\operatorname{F}(b-b,R)=0.$ (10.9)
Since $\operatorname{F}(b-b,R)=1$ we can solve for $\operatorname{C}(b,S)$ to reach Equation 10.7. Since $\operatorname{F}(b,R)\neq\alpha\operatorname{F}(b-a,R)$, $\operatorname{C}(b,S)\neq 0$. When $\operatorname{C}(b,S)$ is negative, $b$ is an antirod.
To prove the converse, suppose that $R$ scales by some $\alpha$ at (ba,b)(b-a,b). Use Equations 10.6 and 10.7 to define $\operatorname{C}(a,S)$ and $\operatorname{C}(b,S)$.
Let $Q$ be the rod set that expands $R$ to the $S$ with shape $\langle a,b\rangle$ and those net rod counts. It follows that the discrepancy $\operatorname{C}(n,Q)=\operatorname{D}(n,R,S)=0$ when $b-\max R<n\leq b$. Corollary 27 implies $Q$ is finite. ∎
When a rod set is 1- or 2-expandable there is a unique $S$ for for each achievable shape: the existence theorems above determine the multiplicities in $S$. For $3$-expandability and higher it’s easy to find counter-examples to uniqueness. For example,
$1,2,2]\xrightarrow{\hphantom{a}[1,2]\hphantom{a}}[2^{2},3^{3},4^{2}]$
and
$1,2,2]\xrightarrow{\hphantom{a}[1,2,2]\hphantom{a}}[2,3^{4},4^{4}]$
have the same shape, $\langle 2,3,4\rangle$.
11 Fibonacci and Lucas 2-Expansions
In this section we study a family of rod sets with shape $\langle 1,2\rangle$ that generalize the Fibonaccis. Their 2-expansions establish divisibility properties of their net train count sequences.
Definition 53.
Rod set $R$ is a Lucas rod set when $R=[1^{s},2^{t}]$ or ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}}^{s},2^{t}]$ for relatively prime positive integers $s$ and $t$. The net train counts for a Lucas rod set form a Lucas sequence.
Lucas rod sets correspond to recursions of the form
$\operatorname{F}(n)=\pm s\operatorname{F}(n-1)+t\operatorname{F}(n-2)$
for positive relatively prime integers $s$ and $t$, with our usual initial conditions.
Be warned that the traditional Lucas numbers, discussed at Equation 5.5, that satisfy the Fibonacci recursion with different initial conditions, are not a Lucas sequence in our sense. In the literature on Lucas sequences, for example [1], [11], and [12], the definitions are different. Ours are right for our rod set model.
Lemma 54.
Let $R$ be a Lucas rod set with $s>1$. Then for $n\geq 1$, $s\mid\operatorname{F}(n,R)$ if and only if $n$ is odd.
Proof.
When $R$ is $1^{s},2^{t}]$ there are $s$ colors for rods of length $1$, so for $n\geq 1$, $s$ divides the number of trains of length $n$ that contain a rod of length $1$. If $n$ is odd, all trains of length $n$ must include at least one rod of length $1$.
When $n\geq 2$ is even, the only trains with no rod of length $1$ are those made of all $2$s; there are $t^{n/2}$ of them. Since $s$ divides the number of rods that contain a rod of length $1$, whether or not it divides $\operatorname{F}(n)$ depends on whether or not it divides $t^{n/2}$. Since by definition, $\gcd(s,t)=1$, $s$ does not divide $t^{n/2}$.
Since odd sign swap 20 preserves the magnitudes of the net train counts, this argument works for ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}}^{s},2^{t}]$ as well. ∎
The next theorem systematically constructs three classes of expansions of Lucas rod sets to a rod set with a two element shape.
Theorem 55.
For each of the following pairs $a<b$ of positive integers, the Lucas rod set $R$ expands via a finite rod set to a unique rod set with shape $\langle a,b\rangle$.
  1. 1.
    $(a,a+1)$.
  2. 2.
    $(a,a+2)$, when $s=1$ or when $a$ is even.
  3. 3.
    $(kd,(k+1)d)$ for $d>2$ and $k\geq 1$.
Proof.
We need prove only existence; Theorem 52 guarantees uniqueness.
It suffices to find the expansions for the positive Lucas rod sets $1^{s},2^{t}]$. Then the invariance of the absolute values of train counts under odd sign swap 20 implies the expansions exist for ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}}^{s},2^{t}]$.
We will illustrate the proof using the running example $1^{3},2^{2}]$, which has net train counts starting at $\operatorname{F}(0)=1$
$1,3,11,39,139,495,1763,6279,22363,79647,283667\ldots.$ (11.1)
Every other count is a multiple of $3$, as predicted by Lemma 54.
In a few cases below when the abstract proof is messy and adds no additional insight, we argue only with the example.
  1. 1.
    To find the expansion to shape $\langle a,a+1\rangle$ we start from $1^{s},2^{t}]=[1^{\operatorname{F}(1)},2^{t\operatorname{F}(0)}]$ and repeatedly expand all of the smallest rods. In the example, we have
    $1^{3},2^{2}]\xrightarrow{\hphantom{a}\hphantom{a}}[2^{11},3^{6}]\xrightarrow{\hphantom{a}\hphantom{a}}[3^{39},4^{22}]\xrightarrow{\hphantom{a}\hphantom{a}}[4^{139},5^{78}]\xrightarrow{\hphantom{a}\hphantom{a}}\ldots.$ (11.2)
    The algebra that computes the exponent of the smaller rod is the same algebra that computes the next term in the recursion. At each step the expanded rod set is
    $a^{\operatorname{F}(a)},(a+1)^{t\operatorname{F}(a-1)}].$
  2. 2.
    To find the expansion to shape $\langle a,a+2\rangle$ when $s=1$ or $a$ is even, we note that Lemma 54 implies those conditions hold just when $t\operatorname{F}(a-1)/s$ is an integer. Then we start with any rod set in Equation 11.2 and add $t\operatorname{F}(a-1)/s$ rod/antirod pairs of length $a$, expand the new antirods, and simplify. For example
    $4^{139},5^{78}]\equiv[4^{139},4^{26},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}}^{26},5^{78}]\xrightarrow{\hphantom{a}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}}^{26}]\hphantom{a}}[4^{165},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{5}}^{78},5^{78},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{6}}^{52}]\equiv[4^{165},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{6}}^{52}].$
  3. 3.
    To find the expansion to shape $\langle kd,(k+1)d\rangle$ for $d>2$ and $k\geq 1$ we first explain how to find examples of shape $\langle d,2d\rangle$.
    We start with the rod set $d^{\operatorname{F}(d)},(d+1)^{t\operatorname{F}(d-1)}]$ in Equation 11.2 and add $t\operatorname{F}(d-2)$ rod/antirod pairs of length $d$. We begin with the antirods, then repeatedly expand the middle length rods until we just have two lengths. Below are the expansions for $d=4$ in our example,
    $[4^{139},5^{78}]$ $\equiv[4^{139},4^{22},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}}^{22},5^{78}]$
    $\xrightarrow{\hphantom{a}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}}^{22}]\hphantom{a}}[4^{161},5^{12},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{6}}^{44}]$
    $\xrightarrow{\hphantom{a}[5^{12}]\hphantom{a}}[4^{161},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{6}}^{8},7^{24}]$
    $\xrightarrow{\hphantom{a}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{6}}^{8}]\hphantom{a}}[4^{161},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{8}}^{16}].$
    The general case provides insight that may not be evident in the example. We let $k=\operatorname{F}(d)+t\operatorname{F}(d-2)$ and omit intermediate steps,
    $[d^{\operatorname{F}(d)},(d+1)^{t\operatorname{F}(d-1)}]\xrightarrow{\hphantom{a}\hphantom{a}}$ $[d^{k},(d+1)^{t^{2}\operatorname{F}(d-2)},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{(d+2)}}^{t^{2}\operatorname{F}(d-1)}]\xrightarrow{\hphantom{a}\hphantom{a}}$
    $[d^{k},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{(d+2)}}^{t^{3}\operatorname{F}(d-3)},(d+3)^{t^{3}\operatorname{F}(d-2)}]\xrightarrow{\hphantom{a}\hphantom{a}}\ldots$
    At the $d$th expansion we reach the rod set $d^{k},(2d)^{t^{d}}],$ where the larger rods are antirods just when $d$ is even.
    To find more expansions with shape lengths that differ by $d=4$, we start with the rod set $4^{161},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{8}}^{16}]$ and repeatedly expand the smallest rods. Corollary 22 implies the expansions we create are also expansions of $1^{3},2^{2}]$ by finite $Q$. Below are the first few such expansions in our example:
    $[4^{161},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{8}}^{16}]$ $\xrightarrow{\hphantom{a}[4^{161}]\hphantom{a}}[8^{(161)(161)-16},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{12}}^{(161)(16)}]$
    $=\phantom{aaaaa}[8^{25905},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{12}}^{2576}]$
    $\xrightarrow{\hphantom{a}[8^{25905}]\hphantom{a}}[12^{(25905)(161)-2576},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{16}}^{(2576)(16)}]\xrightarrow{\hphantom{a}\hphantom{a}}\ldots$ (11.3)
    These show we have an expansion to shape $\langle 4k,4(k+1)\rangle$ for all $k$.
To motivate the next theorem, which proves Lucas sequences are divisibility sequences, we look again at the train counts in 11.1, where we know that every other count is divisible by $\operatorname{F}(1)=3$. Much more is true. Every third count ($11,495,22363\ldots$) is divisible by $\operatorname{F}(2)=11$, every fourth count ($39,6279,\ldots$) by $\operatorname{F}(3)=39$, and so on.
Definition 56.
An integer sequence $\mathscr{L}$ starting at index 1 is a divisibility sequence when
$m\mid n\implies\mathscr{L}(m)\mid\mathscr{L}(n).$
Theorem 57.
The net train counts for a Lucas rod set $R$, reindexed so that $\mathscr{L}(n)=\operatorname{F}(n-1,R)$, form a divisibility sequence.
Proof.
Since odd sign swaps preserve the magnitude of the train counts, we can assume $R=[1^{s},2^{t}]$. We need to show that
$(m-1)\mid(n-1)\implies\operatorname{F}(m-1)\mid\operatorname{F}(n-1).$ (11.4)
Theorem 55 proves the existence of the expansions, which together with Equation 10.4 in the 2-shape Theorem 52 imply the needed divisibility relationships between net train counts:
  1. 1.
    The expansions to rod sets with lengths $\langle a,a+2\rangle$ for odd $a$ imply the divisibilities $\operatorname{F}(1)=s\mid\operatorname{F}(a)$ for odd $a$. This is not a surprise, since we needed Lemma 54 to find those expansions.
  2. 2.
    For $d>2$, the existence of expansions to rod sets with lengths $\langle kd,(k+1)d\rangle$ implies $\operatorname{F}(d-1)\mid\operatorname{F}(kd-1)$.
  3. 3.
    For the last kind of divisibility we return to our example. Equation 3 shows how to expand $R=[1^{3},2^{2}]$ with finite $Q$ to $4^{161},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{8}}^{16}]$, to $8^{25905},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{12}}^{2576}]$, and so on. The second of these expansions implies that $R$ scales by $25905$ at (128,12)(12-8,12), which means that $\operatorname{F}(11)=25905\operatorname{F}(3)$ so $\operatorname{F}(3)|\operatorname{F}(11)$.
An integer sequence $\mathscr{L}$ starting at index 1 is a strong divisibility sequence when for all $m,n\geq 1$
$\gcd(\mathscr{L}(m),\mathscr{L}(n))=\mathscr{L}(\gcd(m,n)).$
It’s well known that the traditionally indexed Fibonacci numbers are a strong divisibility sequence. It’s somewhat less well known that the Lucas sequences, as we have defined them are also strong divisibility sequences [1][5][6]. This result implies that the expansions we found in Theorem 55 lead to all possible two-term linear recursions involving Lucas sequences. In particular, we have found all such Fibonacci recursions.
12 Borwein trinomials
[…]
Definition 58.
A Borwein polynomial is a polynomial all of whose coefficients are $\pm 1$. We will assume in addition that the constant term is 1.
Borwein polynomials correspond to finite rod sets with just one rod of each length.
[…]
Definition 59.
Finite rod sets $S$ and $S^{\prime}$ are congruent modulo $n$ when there is a bijection $\pi$ from $S$ to $S^{\prime}$ such that for all $k\in S$, $\pi(k)\equiv k\pmod{n}$ and have the same sign.
Theorem 60.
[…]
In each case the multiplicities of $a$ and $b$ are $\pm 1$.
Proof.
We write down the period $6$ or period $3$ net train counts for each rod set. Since the counts are all $\pm 1$ or 0 Theorem 52 says the expansions correspond to a scaling that matches 1’s or 1-1s at (ba1,b1)(b-a-1,b-1) with $\alpha=1$ or one that matches 1 and 1-1 with $\alpha=-1$. ∎
13 $2$-Expandability when $\max R>2$
So far we have applied Theorem 52 when $R$ contains only rods of lengths 1 and 2. We end this section (and the paper) with some examples and conjectures on $2$-expansions of rod sets $R$ with $\max R>2$.
When R=[1,2,,k]R=[1,2,...,k] the the net train counts for $R$ start with consecutive powers of $2$ and scale for a while at the start. A finite variant of Equation 7.3 leads to the expansion
$1,2,...,k]\xrightarrow{\hphantom{a}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}}]\hphantom{a}}[1^{2},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k+1}}].$
Thus for any finite $k$, there exists at least one $2$-expandable rod set $R$ with $\max R=k$.
Nonetheless, when $\max R\geq 3$ the scalability conditions are quite restrictive. There are just three rod sets with two different positive rods of length at most $3$. We’ve studied the Fibonacci rod set $[1,2]$. To find 2-expansions of the Padovan rod set, $[2,3]$, we examine the net train count sequence
$P(n)=1,0,1,1,1,2,2,3,4,5,7,9,12,16,21,28,37,49,86,114,151,200,\ldots$
looking for scalings: two pairs of consecutive terms (P(ba2),P(ba1))(P(b-a-2),P(b-a-1)) and (P(b2),P(b1))(P(b-2),P(b-1)), where the second pair is an integer multiple of the first. This happens with the net train counts $1,1,1$ when two $(1,1)$’s overlap, which corresponds to the expansion to $[1,5]$.
Then twice we see pairs of consecutive net train counts $(1,1)$ and $(2,2)$. Further along we find $(3,4)$ and its multiples $(9,12),(12,16)$, and $(21,28)$. These lead to the following Padovan expansions (and also to their odd sign swaps):
$[2,3]$ $\xrightarrow{\hphantom{a}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}},2]\hphantom{a}}[1,5]$
$[2,3]$ $\xrightarrow{\hphantom{a}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}},3,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}}]\hphantom{a}}[2,2,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{7}}]=[2^{2},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{7}}]$
$[2,3]$ $\xrightarrow{\hphantom{a}[2,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}},4]\hphantom{a}}[3^{2},7]$
$[2,3]$ $\xrightarrow{\hphantom{a}\hphantom{a}}[4^{3},13]$
$[2,3]$ $\xrightarrow{\hphantom{a}\hphantom{a}}[5^{4},14]$
$[2,3]$ $\xrightarrow{\hphantom{a}\hphantom{a}}[7^{7},16^{2}].$
Those are all the scalings we could see simply staring at the sequence. Note that past $\operatorname{F}(15)=28$ the numbers seem to be growing in a way that makes further such pairs unlikely, which we confirmed with a computer search up to $n=1000$.
Similar investigations by hand and by computer up to $n=1000$ yield three finite expansions of the Narayana rod set $[1,3]$, to $2^{2},7]$, $3^{3},8]$, and $11^{67},33]$.
When we allow antirods but still require $\max R=3$ and one kind of each rod we find the following expansions (and their odd sign swaps).
[…]
There are $2$-expansions from the rod set ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}},3]$ to four more rod sets: ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{5}}^{2},7]$, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{5}}^{3},8]$, $7^{3},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{8}}^{2}]$, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{22}}^{67},33]$.
We’ve found only two $2$-expansions with $\max R>3$:
$1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}}]\xrightarrow{\hphantom{a}[1,2,3,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{5}},6,9]\hphantom{a}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{6}}^{3},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{13}}]$ (13.1)
and
$3,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}}]\xrightarrow{\hphantom{a}[3,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}},6,7,8,9]\hphantom{a}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{7}}^{3},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{13}}]$ (13.2)
What next?
  • There is more to discover on the connections between expansions of $R$ and the structure of the net train count sequence $\operatorname{F}(n,R)$. Theorems 50 and 52 settle those connections for expansions to shapes of cardinality 1 and 2. What are the corresponding theorems for larger shapes?
  • When finite rod set $R$ contains just one rod of length $m=\max R$ it has a natural reflection: the rod set whose rod set polynomial is $x^{m}\operatorname{\mathcal{C}}(1/x,R)$. For example, Equations 13.1 and 13.2 are reflections of each other, as are other examples in the last section—the Narayana expansion $11^{67},33]$ and ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{22}}^{67},33]$, which is an expansion of ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}},3]$. When a rod set has a reflection we can use its recursion to define net train counts for trains with “negative length”. What are the theorems that relate reflection to expansion and to net train counts?
  • Much of our rod set algebra works when rod lengths are elements of an ordered group. We use the fact that lengths are positive integers only as exponents in the rod generating functions and when discussing the recursions determined by rod sets. Is any of this generalization useful or surprising?
  • Our theorems on finite expansions translate to theorems about polynomial multiplication. What we don’t have are theorems that say there are no finite expansions to some particular rod sets. Such theorems would be rod set algebra proofs of irreducibility. Are there any?
  • There is substantial interest in factoring polynomials over finite fields. Can our rod and tree model for polynomial arithmetic be of use?
  • We have completely avoided analysis. In particular, we discovered some of early examples of expansion by looking at the growth rates of train counts. Can our rod set model say anything useful about train count growth rates?

### v2

@@
 We complement abstractions with lots of examples.
+keywords:
+recursion, compositions, Fibonacci, Padovan, generating function, Lucas sequence, Borwein polynomial
 Introduction
@@
 Soon we will use our rod model to study this observation about the Fibonacci sequence by converting trains composed of rods of lengths $1$ and $2$ into trains composed of rods of lengths $1$, 33, and $4$.
-When a 22 is followed by a $1$ we can glue them together to make a $3$, and when 22 is followed by a $2$ we glue to make a $4$.
+When a $2$ is followed by a $1$ we can glue them together to make a $3$, and when $2$ is followed by a $2$ we glue to make a $4$.
 This construction may end with an extra rod of length $2$ at the end of the train.
@@
 We start with careful definitions and some elementary observations about rods, rod sets and rod trains.
-Definition 1.
+Definition 1.1.
 A rod is a triple $(r,c,\pm 1)$ in which the length $r$ is a positive integer and $c$ is an arbitrary tag we think of as a color.
@@
 Formally, if $r=(m,c,\pm 1)$ then nr=nmn-r=n-m.
-Definition 2.
+Definition 1.2.
 A rod train, or simply a train, is a finite sequence of rods.
@@
 For example, $|{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}}|=|1{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}1|=|11{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}|=|1.\Lambda.1{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}|={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}}$.
-Definition 3.
+Definition 1.3.
 A rod set is a set of rods with finitely many rods of each length.
@@
 $2,2,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}},3]\setminus[2]=[2,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}},3].$
-Definition 4.
+Definition 1.4.
 A rod train set, or simply a train set, is a set of rod trains. The net train count $\operatorname{C}(n,T)$ is the difference between the number of positive and negative trains of that length.
@@
 $|\operatorname{Trains}([\phantom{a}])|$ is empty because $\operatorname{Trains}([\phantom{a}])=\{\Lambda\}$ and there is no rod $|\Lambda|$.
-Definition 5.
+Definition 1.5.
 For rod set $R$, the net train count sequence is
@@
 Table 1: Rod trains for $1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}]$
-Theorem 6.
+Theorem 1.6.
 For rod set $R$, the net train count sequence $\operatorname{F}(n,R)$ satisfies the recursion
@@
   • $\operatorname{F}(n,[1,1])=1,1,2,4,8,\ldots$, the powers of $2$.
-  • $\operatorname{F}(n,[2,3])=1,0,1,1,1,2,2,3,4,5,7,9,12,\ldots$, a Padovan sequence [14].
+  • $\operatorname{F}(n,[2,3])=1,0,1,1,1,2,2,3,4,5,7,9,12,\ldots$, a Padovan sequence [8, p 89] [16].
   • $\operatorname{F}(n,[1,3])=1,1,1,1,2,3,4,6,9,13,19,\ldots$, a Narayana cow number sequence [15].
@@
 The next theorem shows that studying the net train count sequences of rod sets is quite general.
-Theorem 7.
+Theorem 1.7.
 For every infinite sequence of integers $a_{1},a_{2},\ldots$ there is a rod set $R$ such that $\operatorname{F}(n,R)=a_{n}$ for all $n\geq 1$.
@@
 $R=[1,2,3^{2},4^{5},5^{14},6^{42},7^{132},8^{429},\ldots].$
-The number of rods of each length is the Catalan number with index shifted by $1$ from the train counts. The recursion corresponds to the well-known convolution that the Catalan’s satisfy.
+The number of rods of each length is the Catalan number with index shifted by $1$ from the train counts. The recursion corresponds to the well-known convolution that the Catalans satisfy.
 2 Trees
-Here we explore a natural way to organize $\operatorname{Trains}(R)$ in a tree111 Sorry for the awkward metaphor mixing trees and trains.
+Here we explore a natural way to organize $\operatorname{Trains}(R)$ in a tree [^1] Sorry for the awkward metaphor mixing trees and trains.
 that lets us sytematically study the way we combined rods in Figure 2.
-Definition 8.
+Definition 2.1.
 Let $R$ be a rod set.
@@
 The next definition formalizes this kind of expansion.
-Definition 9.
+Definition 2.2.
 A rooted subtree $\mathfrak{T}$ of $\operatorname{Tree}(R)$ is an expansion subtree when
@@
 We begin with a definition that refers only to rod sets and recursions, not to subtrees.
-Definition 10.
+Definition 3.1.
 For rod sets $R$ and $S$ the discrepancy at $n>0$ is the difference
@@
 which will follow from the next theorem.
-Lemma 11.
+Lemma 3.2.
 (Parsing) Suppose $\mathfrak{T}$ is an expansion subtree of $\operatorname{Tree}(R)$. Then we can write every nonempty $R$-train $\rho$ uniquely as a concatenation
@@
 […]
-Theorem 12.
+Theorem 3.3.
 (Discrepancies) Suppose $\mathfrak{T}$ is an expansion subtree of $\operatorname{Tree}(R)$.
@@
 Proof.
-Let $A$ be the subset of $\operatorname{Trains}(R)$ for which $m=0$ in Lemma 11, and let $B=\operatorname{Trains}(R)\setminus A$.
+Let $A$ be the subset of $\operatorname{Trains}(R)$ for which $m=0$ in Lemma 3.2, and let $B=\operatorname{Trains}(R)\setminus A$.
 For trains of length $n$, we have:
@@
 The next theorem shows that this kind of equality occurs only in the obvious way.
-Theorem 13.
+Theorem 4.1.
 The following assertions about rod sets $R$ and $S$ are equivalent.
@@
 The proof is an elementary consequence of the definitions; we leave it to the reader. ∎
-Definition 14.
+Definition 4.2.
 Rod sets $R$ and $S$ are equivalent, written $R\equiv S$, when they satisfy any one (and hence all) of the conditions in the previous theorem.
@@
 When $R$ is finite we write $\|R\|$ for the number of rods in $T(R)$ and $\max R$ for the length of the longest rod in $T(R)$.
-Theorem 15.
+Theorem 4.3.
 If $R\equiv R^{\prime}$ and $S\equiv S^{\prime}$ then
@@
 ∎
-Corollary 16.
+Corollary 4.4.
 The rod set $|\operatorname{Trains}(R)|$ is equivalent to the reduced rod set containing $\operatorname{F}(n,R)$ rods of length $n$ with appropriate signs.
@@
 If all the rods $Q$ we want to expand are already in $R$ then the rod set $|\text{Leaves}(\mathfrak{T})|$ for the expansion subtree $\mathfrak{T}$ of $\operatorname{Tree}(R)$ with $|\text{Inner}(\mathfrak{T})|=Q$ is $(R\setminus Q)\cup QR$.
-Theorem 15 then implies
+Theorem 4.3 then implies
 $|\text{Leaves}(\mathfrak{T})|=(R\setminus Q)\cup QR\equiv R\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}\cup QR.$
 We will soon see that up to rod set equivalence this union is the leaf rod set when expanding $R$ by $Q$, even when $Q$ is not a subset of $R$.
-Definition 17.
+Definition 5.1.
 For rod sets $R$, $Q$ and $S$ we say $R$ expands to $S$ via $Q$, and write
@@
 $R^{\prime}\xrightarrow{\hphantom{a}\text{Inner}(\mathfrak{T})\hphantom{a}}\text{Leaves}(\mathfrak{T}).$
-Theorem 18.
+Theorem 5.2.
 (Expansion) For rod sets $R$ and $Q$ there is a rod set $S$ such that
@@
 Now we explore some simple consequences of this theorem that define useful operations in our rod set algebra.
-Corollary 19.
+Corollary 5.3.
 (Exchange) Rod set $R$ expands via $Q$ to $S$ if and only if $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}$} expands via $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{R}$} to $S$.
@@
 $R\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}\cup QR={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}\cup R\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}.{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{R}}.$
-Corollary 20.
+Corollary 5.4.
 (Odd Sign Swap) Suppose $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$.
@@
 The next theorem collects several important equivalent ways to characterize the expansion $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$.
-Theorem 21.
+Theorem 5.5.
 Let $R$, $Q$ and $S$ be rod sets. Then the following assertions are equivalent.
@@
 Proof.
-We use any convenient expansion subtree $\operatorname{Tree}(R^{\prime})$ for a rod set equivalent to $R$ and invoke Lemma 11 and Theorem 12. ∎
+We use any convenient expansion subtree $\operatorname{Tree}(R^{\prime})$ for a rod set equivalent to $R$ and invoke Lemma 3.2 and Theorem 3.3. ∎
 As a corollary we see how to compose expansions:
-Corollary 22.
+Corollary 5.6.
 (Composition) Let $P,R,S$ be rod sets with $P\xrightarrow{\hphantom{a}Q_{PR}\hphantom{a}}R$,
@@
 We distribute, view $Q_{RS}.Q_{PR}$ as one suffix, and the corollary follows. ∎
-Theorem 18 shows that there is a simple formula for $S$ in terms of $R$ and $Q$. The next theorem proves that given any two of $R,Q,S$, the third is well-defined.
-Theorem 23.
+Theorem 5.2 shows that there is a simple formula for $S$ in terms of $R$ and $Q$. The next theorem proves that given any two of $R,Q,S$, the third is well-defined.
+Theorem 5.7.
 Any two of $R,Q,S$ uniquely determine the third so that
@@
 Proof.
-Theorem 18 tells us how to find $S$ given $R$ and $Q$.
-Given $R$ and $S$ we can use Theorem 21 to find $Q$ by computing the discrepancies $\operatorname{D}(n,R,S)$.
+Theorem 5.2 tells us how to find $S$ given $R$ and $Q$.
+Given $R$ and $S$ we can use Theorem 5.5 to find $Q$ by computing the discrepancies $\operatorname{D}(n,R,S)$.
 […]
@@
 Below we formalize this algorithm. Follow along in Figure 6.
-Algorithm 24.
+Algorithm 5.8.
 Let $R$ and $S$ be reduced rod sets with finite $S$.
@@
 We have been careful to ensure that all our algebra works with infinite rod sets. The next theorems show some of what we can prove when we impose finiteness conditions.
-Theorem 25.
+Theorem 5.9.
 If $Q$ is finite and $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$ then for $n>\max Q$, $\operatorname{Trains}(R)$ satisfies the recursion determined by $S$.
@@
 For finite $R$ and $S$ the $Q$ in $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$ need not be finite.
-For example, when we use Algorithm 24 to find the $Q$ that mediates the expansion of the Fibonacci rod set $[1,2]$ to the Padovan rod set $[2,3]$ we expand the $1$, add a $3/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}}$ pair, then expand the $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}$} forever.
+For example, when we use Algorithm 5.8 to find the $Q$ that mediates the expansion of the Fibonacci rod set $[1,2]$ to the Padovan rod set $[2,3]$ we expand the $1$, add a $3/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}}$ pair, then expand the $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}$} forever.
 The result is
@@
 The next theorem generalizes this observation.
-Theorem 26.
+Theorem 5.10.
 Suppose $R$ and $S$ are finite rod sets and $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$.
@@
 $\operatorname{C}(n,Q)=\sum_{k\in U}\operatorname{F}(n-k,R).$
-Corollary 27.
+Corollary 5.11.
 If $R$ and $S$ are finite rod sets with $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$, then $Q$ is finite if and only if starting for some $N\geq\max S-\max R$ there are $\max R$ consecutive rod lengths $k$ for which $\operatorname{C}(k,Q)=0$.
@@
 $1,\mathscr{L}=1,2,1,3,4,7,11,18,\ldots,$ (5.7)
-then Theorem 7 implies there is a rod set $R$ whose net train counts are $1,\mathscr{L}$.
+then Theorem 1.7 implies there is a rod set $R$ whose net train counts are $1,\mathscr{L}$.
 But that rod set is uninformative: it does not clearly connect the Lucas number sequence to the Fibonacci recursion determined by the rod set $[1,2]$.
-However, Theorem 21 says that the rod set $Q$ for which
+However, Theorem 5.5 says that the rod set $Q$ for which
 $R\xrightarrow{\hphantom{a}Q\hphantom{a}}[1,2]$
@@
 and $Q=[1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}]$.
-Thus, also by Theorem 21, the Lucas numbers satisfy the recursion
+Thus, also by Theorem 5.5, the Lucas numbers satisfy the recursion
 $\mathscr{L}(n)=\operatorname{F}(n,[1,2])+\operatorname{F}(n-1,[1,2])-2\operatorname{F}(n-2,[1,2]),$ (5.8)
 which expresses each Lucas number in terms of Fibonacci numbers with default initial conditions.
-Theorem 21 provides more. Equation 5.3 shows a nice way to view the trains counted by $\mathscr{L}$:
+Theorem 5.5 provides more. Equation 5.3 shows a nice way to view the trains counted by $\mathscr{L}$:
 $|\operatorname{Trains}(R)|\equiv|\operatorname{Trains}([1,2])|\cup|\operatorname{Trains}([1,2]).[1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}]|.$
@@
 Table 2: Rod trains for the Lucas numbers
-The argument in this Lucas number example relies only on Theorem 21, not on the finiteness of $S=[1,2]$ nor on the fact that the discrepancies occur only for values $n\leq\max S$.
+The argument in this Lucas number example relies only on Theorem 5.5, not on the finiteness of $S=[1,2]$ nor on the fact that the discrepancies occur only for values $n\leq\max S$.
 But when these finiteness conditions are met the result shows how to use our rod model to define “traditional” initial conditions.
@@
 Our algebraic operations on rod sets define interactions between two different formal power series: the rod count generating function and the recursion generating function.
-Definition 28.
+Definition 6.1.
 The rod count generating function for rod set $R$ is the formal power series
@@
 The next lemma establishes basic properties of the rod count generating function.
-Lemma 29.
+Lemma 6.2.
 For rod sets $Q,R$
@@
 We add or subtract this quantity from $1$ as needed, as illustrated in the next theorem.
-Theorem 30.
+Theorem 6.3.
 Rod set $R$ expands to $S$ via $Q$ if and only if
@@
 Proof.
-If $R$ expands to $S$ via $Q$ then Theorem 18 implies $S\equiv R\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}\cup QR$ so
+If $R$ expands to $S$ via $Q$ then Theorem 5.2 implies $S\equiv R\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}\cup QR$ so
 $1-\operatorname{\mathcal{C}}(x,S)$ $=1-(\operatorname{\mathcal{C}}(x,R)-\operatorname{\mathcal{C}}(x,Q)+\operatorname{\mathcal{C}}(x,QR))$
@@
 $(1-x-x^{3}-x^{5}-\cdots)=(1-x-x^{2})(1+x^{2}+x^{4}+\cdots).$
-The Exchange Corollary 19 is an immediate consequence of Theorem 30. The Odd Sign Swap Corollary 20 follows when we replace $x$ with x-x in Equation 6.3.
-This theorem shows that the Expansion Theorem 18 is essentially equivalent to the standard algorithm for multiplying formal power series.
+The Exchange Corollary 5.3 is an immediate consequence of Theorem 6.3. The Odd Sign Swap Corollary 5.4 follows when we replace $x$ with x-x in Equation 6.3.
+This theorem shows that the Expansion Theorem 5.2 is essentially equivalent to the standard algorithm for multiplying formal power series.
 The net count of the rods of length $n$ in $R\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}\cup QR$ encodes the calculation of the coefficient of $x^{n}$ in the product power series.
-Then the Composition Corollary 22 is just the distributive property.
-Algorithm 24 illustrated in Figure 6 calculates the quotient of two power series.
+Then the Composition Corollary 5.6 is just the distributive property.
+Algorithm 5.8 illustrated in Figure 6 calculates the quotient of two power series.
 While working on this paper we often found that we made fewer arithmetic errors when we calculated expansions using trees rather than by multiplying and dividing polynomials or power series using the standard algorithms.
 However, we do not recommend teaching these methods in schools.
-Corollary 31.
+Corollary 6.4.
 If $R$ is a finite rod set then the rod count generating functions for the rod sets $S$ that are expansions of $R$ by a finite $Q$ are precisely the polynomials with constant term $1$ in the ideal generated by $1-\operatorname{\mathcal{C}}(x,R)$ in the polynomial ring $\mathbb{Z}[x]$.
@@
 then $q(x)$ has constant term $1$ and $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$ for the rod sets $S$ and $Q$ corresponding to the polynomials $s(x)$ and $q(x)$. ∎
-The next corollary is a reprise of Theorem 23.
-Corollary 32.
+The next corollary is a reprise of Theorem 5.7.
+Corollary 6.5.
 Any two of $R,Q,S$ uniquely determine the third so that $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$.
@@
 Using Equation 6.3 we can solve for any one of the rod count generating functions in terms of the other two. ∎
-Theorem 33.
+Theorem 6.6.
 For rod set $R$
@@
 $=(1-\operatorname{\mathcal{C}}(x,R))\operatorname{\mathcal{F}}(x,R).$
-The second assertion follows from Theorem 30. ∎
+The second assertion follows from Theorem 6.3. ∎
 For example, the recursion generating function for the Fibonaccis is
@@
 which nicely mirrors the decomposition of $\operatorname{Trains}(R)$ in Equation 1.1.
-The next corollary is a reprise of Theorem 26.
-Corollary 34.
+The next corollary is a reprise of Theorem 5.10.
+Corollary 6.7.
 Suppose $R$ and $S$ are finite rod sets and $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$. Then for $n>\max S$, $\operatorname{C}(n,Q)$ satisfies the recursion determined by $R$.
@@
 ∎
-The following theorem is a reprise of Theorem 7.
-Theorem 35.
+The following theorem is a reprise of Theorem 1.7.
+Theorem 6.8.
 Every formal power series with constant term $1$ is the recursion generating function $\operatorname{\mathcal{F}}(x,R)$ for a rod set $R$ unique up to rod/antirod equivalence.
@@
 Then we let $R$ be the rod set with $b_{k}$ rods (or antirods) of length $k$ . ∎
-We end this section with a well-known theorem on rational generating functions[7]:
-Theorem 36.
+We end this section with a well-known theorem on rational generating functions[17, p 202]:
+Theorem 6.9.
 The net train counts for a rod set $R$ eventually satisfy the recursion for a finite rod set $S$ if and only if $\operatorname{\mathcal{F}}(x,R)$ is a quotient of polynomials.
@@
 $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S.$
-Since $Q$ specifies the discrepancies in this expansion, Theorem 25 implies it is finite, so
+Since $Q$ specifies the discrepancies in this expansion, Theorem 5.9 implies it is finite, so
 $\operatorname{\mathcal{F}}(x,R)=\frac{1+\operatorname{\mathcal{C}}(x,Q)}{1-\operatorname{\mathcal{C}}(x,S)}$ (6.6)
@@
 Duality is especially useful for exploring relationships between combinatorial identities and compositions, and we finish by deriving some Fibonacci and Padovan examples.
-Definition 37.
+Definition 7.1.
 The dual of rod set $Q$ is
 $Q^{*}=|\operatorname{Trains}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}})|.$
-Theorem 38.
+Theorem 7.2.
 If
@@
 The leaves on the left are ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}^{k}R$ starting at $k=0$, while the leaves on the right are ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}^{k-1}QR$, starting at $k=1$.
-Equation 15 shows that all but the leaves in $R$ cancel.
+Equation 4.3 shows that all but the leaves in $R$ cancel.
 ∎
-Corollary 39.
+Corollary 7.3.
 For rod set $Q$,
@@
 Proof.
-If $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$, then $S\xrightarrow{\hphantom{a}Q^{*}\hphantom{a}}R$ so $R\xrightarrow{\hphantom{a}Q^{**}\hphantom{a}}S$. Then Theorem 18 implies $Q^{**}\equiv Q$. ∎
+If $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$, then $S\xrightarrow{\hphantom{a}Q^{*}\hphantom{a}}R$ so $R\xrightarrow{\hphantom{a}Q^{**}\hphantom{a}}S$. Then Theorem 5.2 implies $Q^{**}\equiv Q$. ∎
 It is easy to write the rod generating function for the dual $Q^{*}$ in terms of the rod generating function for $Q$.
-Theorem 40.
+Theorem 7.4.
 For rod set $Q$, we have
@@
 $1+\operatorname{\mathcal{C}}(x,Q^{*})=\operatorname{\mathcal{F}}(x,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}})=\frac{1}{1-\operatorname{\mathcal{C}}(x,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}})}=\frac{1}{1+\operatorname{\mathcal{C}}(x,Q)}.$
-When $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$, Theorem 18 shows how to compute $S$ given $R$ and $Q$, and Theorem 23 tells us that any two of $R,Q,S$ determines the third.
+When $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$, Theorem 5.2 shows how to compute $S$ given $R$ and $Q$, and Theorem 5.7 tells us that any two of $R,Q,S$ determines the third.
 With duality, it is straightforward, albeit a bit messy, to find explicit rod set formulas for $Q$ in terms of $R$ and $S$ and for $R$ in terms of $Q$ and $S$.
@@
 Then duality reveals an interesting connection between identities involving sums and those involving compositions.
-For example, recall that we established the Fibonacci identities in Equations 3.2 and 3.3 using the third expansion tree in Figure 4 together with Theorem 21. We started with the infinite expansion
+For example, recall that we established the Fibonacci identities in Equations 3.2 and 3.3 using the third expansion tree in Figure 4 together with Theorem 5.5. We started with the infinite expansion
 $1,2]\xrightarrow{\hphantom{a}|\operatorname{Trains}(2)|\hphantom{a}}[1,3,5,7,\ldots],$ (7.1)
@@
 $1+3+8+21+\cdots+\operatorname{F}(2n-1)$ $=\operatorname{F}(2n)-1.$
-Now we use Theorem 38 to dualize Equation 7.1:
+Now we use Theorem 7.2 to dualize Equation 7.1:
 $1,3,5,7,\ldots]\xrightarrow{\hphantom{a}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}]\hphantom{a}}[1,2].$ (7.2)
-Theorem 12 implies the train counts for the rod set $1,3,5,7,\ldots]$ satisfy the Fibonacci recursion when $n>2$.
+Theorem 3.3 implies the train counts for the rod set $1,3,5,7,\ldots]$ satisfy the Fibonacci recursion when $n>2$.
 The train counts for the rod set of odd numbers are the number of compositions, $\mathscr{O}(n)$, of $n$ into odd parts.
@@
 $R=[2,2,3]\xrightarrow{\hphantom{a}|\operatorname{Trains}([3])|\hphantom{a}}[2,2,5,5,8,8,11,11,\ldots].$
-Since $R=[2,2,3]$ is an expansion of the Fibonacci rod set $[1,2]$ by the singleton rod set $[1]$, Corollary 22 implies $1,2]\xrightarrow{\hphantom{a}\hphantom{a}}S$ has discrepancies $3,6,9,\ldots]\cup[1,4,7,\ldots]$.
+Since $R=[2,2,3]$ is an expansion of the Fibonacci rod set $[1,2]$ by the singleton rod set $[1]$, Corollary 5.6 implies $1,2]\xrightarrow{\hphantom{a}\hphantom{a}}S$ has discrepancies $3,6,9,\ldots]\cup[1,4,7,\ldots]$.
 The discrepancies are the rods with lengths $n\equiv 0,1\pmod{3}$, which leads to the following identities for the Fibonacci sequence $\operatorname{F}(n)=\operatorname{F}(n,[1,2])$:
@@
 Even more generally, when $R$ is positive and finite we can (but won’t) systematically compute $\operatorname{F}(n,R)$ as a sum of multinomial coefficients on hyperplanes in high dimensional analogues of the triangle.
-Definition 41.
+Definition 8.1.
 Let $R$ be a rod set and $\tau\in\operatorname{Trains}(R)$. The underlying partition of $\tau$ is the multiset of rods in $\tau$.
 For example, if $R=[3,4]$ and $\tau=34443$ then the underlying partition of $\tau$ is $\{3,3,4,4,4\}$.
-Theorem 42.
+Theorem 8.2.
 Let $R=[a,b]$ be a positive rod set. Then
@@
 The first sum counts rods of length $b$ and the second rods of length $a$.
-Above the dark diagonal, the diagonal for $\operatorname{F}(n,[a,b])=0$ is empty (dashed in the figure) unless $n$ is a multiple of $a$ or $b$.
+Above the dark diagonal, the diagonal for $\operatorname{F}(n,[a,b])=0$ is empty (dashed in the figure) unless $n$ is a linear combination of $a$ and $b$.
 Below the dark diagonal the train counts are positive.
@@
 Figure 8: Train counts for the Fibonaccis, $[1,2]$, and Padovans, $[2,3]$, as diagonal sums in Pascal’s triangle
-Theorem 42 leads to some nice binomial identities.
+Theorem 8.2 leads to some nice binomial identities.
 The train counts for $R=[1,1]$ are powers of two.
@@
 $1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}]\xrightarrow{\hphantom{a}[1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}}]\hphantom{a}}[6].$ (9.2)
-Theorem 25 says that that when $n>4$, the net train counts $\operatorname{F}(n,[1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}])$ satisfy the recursion $\operatorname{F}(n)=\operatorname{F}(n-6)$ determined by rod set $[6]$.
+Theorem 5.9 says that that when $n>4$, the net train counts $\operatorname{F}(n,[1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}])$ satisfy the recursion $\operatorname{F}(n)=\operatorname{F}(n-6)$ determined by rod set $[6]$.
 This example generalizes easily.
-Definition 43.
+Definition 9.1.
 The net train counts for rod set $R$ are eventually periodic with period $p>0$ when there is an $N\geq 0$ such that
@@
 The following theorem characterizes rod sets with eventually periodic net train counts.
-Theorem 44.
+Theorem 9.2.
 (Periodicity) The net train counts for rod set $R$ are eventually periodic with period $p$ if and only if $R$ expands to $[p]$ via a finite $Q$.
@@
 Proof.
-This is the special case of Theorem 25 when $S$ contains just a single positive rod. ∎
+This is the special case of Theorem 5.9 when $S$ contains just a single positive rod. ∎
 Expansions to a single antirod are related to periodicity.
@@
 We generalize in the following theorem.
-Theorem 45.
+Theorem 9.3.
 If $R$ expands to $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}$}] via $Q$, then $R$ expands to $[2k]$ via $Q\cup[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}}]\cup Q{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}}$, and $\operatorname{F}(n,R)$ is eventually periodic with period $2k$.
 Proof.
-Since $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}$}] expands via $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}$}] to $[2k]$, Corollary 22 implies $R$ expands to $[2k]$ via $Q\cup[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}}]\cup Q{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}}$.
+Since $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}$}] expands via $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}$}] to $[2k]$, Corollary 5.6 implies $R$ expands to $[2k]$ via $Q\cup[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}}]\cup Q{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}}$.
 ∎
@@
 The following characterization of periodic finite rod sets includes a proof that eventually periodic finite rod sets are actually periodic.
-Theorem 46.
+Theorem 9.4.
 The following assertions about finite rod set $R$ are equivalent.
@@
 The last three equivalences involve the rod count and recursion generating functions for $R$.
-Theorem 44 says periodicity is equivalent to $R$ expanding to the one element rod set $[p]$, for which $\operatorname{\mathcal{C}}(x,[p])=x^{p}$.
-When $R$ is finite Theorem 30 then implies (1) and (5) are equivalent.
+Theorem 9.2 says periodicity is equivalent to $R$ expanding to the one element rod set $[p]$, for which $\operatorname{\mathcal{C}}(x,[p])=x^{p}$.
+When $R$ is finite Theorem 6.3 then implies (1) and (5) are equivalent.
 The equivalence of (5) and (6) follows from the fact that the complete factorization of $x^{p}-1$ into irreducible polynomials is
@@
 where $\Phi_{d}(x)$ is the cyclotomic polynomial whose roots are the primitive $d$th roots of unity[5].
-The equivalence of (7) is Theorem 33. ∎
+The equivalence of (7) is Theorem 6.6. ∎
 Repeated cyclotomic factors lead to a rod set that isn’t periodic.
@@
 ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}},3,4,5,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{7}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{8}}]\xrightarrow{\hphantom{a}\hphantom{a}}[30].$
-We encourage the reader to check this result using Algorithm 24. Theorem 45 says we can stop when we’ve expanded to $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{15}$}.
+We encourage the reader to check this result using Algorithm 5.8. Theorem 9.3 says we can stop when we’ve expanded to $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{15}$}.
 In the examples of periodic rod sets that we’ve seen so far, there has been at most one rod of each length, but there are periodic rod sets that contain multiple rods of the same length.
@@
 From now on we will often work with expansions where all rod sets are finite. To save repeatedly typing that:
-Definition 47.
+Definition 10.1.
 The expansion $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$ is a finite expansion when all three rod sets $R$, $Q$ and $S$ are finite.
@@
 As usual, we will write $\operatorname{F}(n)$ for $\operatorname{F}(n,R)$ when there is no potential ambiguity.
-Definition 48.
+Definition 10.2.
 A shape is a set of rod lengths, usually written as an ordered list in angle brackets.
@@
 In this example $R\equiv[2^{3},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}}^{1}]$.
-Definition 49.
+Definition 10.3.
 Let $R$ be a rod set and $H$ a shape.
@@
 $\operatorname{F}(n,R)=1,-2,2,0,-4,8,-8,0,16,-32,32,0,-64,128,-128,0\ldots.$ (10.2)
-Since $\max Q=2$, Theorem 25, tells us that the $\operatorname{D}(n,R,[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}}^{4}])$ must vanish for $n\geq 3$.
+Since $\max Q=2$, Theorem 5.9, tells us that the $\operatorname{D}(n,R,[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}}^{4}])$ must vanish for $n\geq 3$.
 In particular, $\operatorname{F}(3)+4F(3-4)=\operatorname{F}(3)=0$ and $\operatorname{F}(4)+4F(0)=-4+4=0$.
 If $\max S-\max R=\max Q$ were greater, then we’d have more discrepancies that would need to be 0. Similar train count conditions will be the key to our theorems.
-Theorem 50.
+Theorem 10.4.
 There is a finite expansion of rod set $R$ to a rod set $S$ with shape $\langle a\rangle$ if and only if $\operatorname{F}(a-i,R)=0$ for $i=1,2,\ldots,\max R-1$ and $\operatorname{F}(a,R)\neq 0$.
@@
 Suppose $\operatorname{Shape}(S)=\langle a\rangle$ and $Q$ is finite in the expansion $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$.
-Theorem 25 implies $a=\max R+\max Q$, so the discrepancies measuring the ways the counts $\operatorname{F}(n,R)$ fail to satisfy the recursion for $S$ are 0 for $n>\max Q=a-\max R$.
+Theorem 5.9 implies $a=\max R+\max Q$, so the discrepancies measuring the ways the counts $\operatorname{F}(n,R)$ fail to satisfy the recursion for $S$ are 0 for $n>\max Q=a-\max R$.
 Thus for positive $i<\max R$
@@
 Setting $\operatorname{C}(a,S)=\operatorname{F}(a,R)$ guarantees discrepancy 0 at $n=a$.
-Then Corollary 27 implies $Q$ is finite.
+Then Corollary 5.11 implies $Q$ is finite.
 ∎
@@
 Later we will show how to find this and all similar two-term Fibonacci identities.
-In Theorem 50 we saw that 1-expansions of $R$ correspond to sequences of 0’s in $\operatorname{F}(n,R)$.
+In Theorem 10.4 we saw that 1-expansions of $R$ correspond to sequences of 0’s in $\operatorname{F}(n,R)$.
 For 2-expandability the correspondence is to divisibility among net train counts.
@@
 To prove our general theorem on $2$-expandability, we formally define scaling.
-Definition 51.
+Definition 10.5.
 The finite rod set $R$ scales at counts $(c,b)$ when there is a unique integer $\alpha$ such that
@@
 $2,3]\xrightarrow{\hphantom{a}Q=[2,3,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}}^{2},5^{2},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{6}},8,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{9}},10]\hphantom{a}}[4^{3},13].$
-The Exchange Corollary 19 implies
+The Exchange Corollary 5.3 implies
 ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}\xrightarrow{\hphantom{a}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}}]\hphantom{a}}[4^{3},13].$ (10.5)
@@
 The next theorem shows that scaling is the heart of $2$-expandability.
-Theorem 52.
+Theorem 10.6.
 There is a finite expansion of rod set $R$ to a rod set $S$ with shape $\langle a,b\rangle$ if and only if there is a nonzero integer $\alpha$ such that $R$ scales by $\alpha$ at (ba,b)(b-a,b), and $\operatorname{F}(b,R)\neq\alpha\operatorname{F}(b-a,R)$.
@@
 Suppose there is a finite expansion $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$ with $\operatorname{Shape}(S)=\langle a,b\rangle$.
-Theorem 25 implies $b=\max R+\max Q$, so the discrepancies measuring the ways the counts $\operatorname{F}(n,R)$ fail to satisfy the recursion for $S$ are 0 for $n>b-\max R$.
+Theorem 5.9 implies $b=\max R+\max Q$, so the discrepancies measuring the ways the counts $\operatorname{F}(n,R)$ fail to satisfy the recursion for $S$ are 0 for $n>b-\max R$.
 In the range $b-\max R<n<b$, we have $\operatorname{F}(n-b,R)=0$.
@@
 It follows that the discrepancy $\operatorname{C}(n,Q)=\operatorname{D}(n,R,S)=0$ when $b-\max R<n\leq b$.
-Corollary 27 implies $Q$ is finite.
+Corollary 5.11 implies $Q$ is finite.
 ∎
@@
 In this section we study a family of rod sets with shape $\langle 1,2\rangle$ that generalize the Fibonaccis. Their 2-expansions establish divisibility properties of their net train count sequences.
-Definition 53.
+Definition 11.1.
 Rod set $R$ is a Lucas rod set when $R=[1^{s},2^{t}]$ or ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}}^{s},2^{t}]$ for relatively prime positive integers $s$ and $t$.
@@
 Ours are right for our rod set model.
-Lemma 54.
+Lemma 11.2.
 Let $R$ be a Lucas rod set with $s>1$. Then for $n\geq 1$, $s\mid\operatorname{F}(n,R)$ if and only if $n$ is odd.
@@
 Since by definition, $\gcd(s,t)=1$, $s$ does not divide $t^{n/2}$.
-Since odd sign swap 20 preserves the magnitudes of the net train counts, this argument works for ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}}^{s},2^{t}]$ as well. ∎
+Since odd sign swap 5.4 preserves the magnitudes of the net train counts, this argument works for ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}}^{s},2^{t}]$ as well.
+∎
 The next theorem systematically constructs three classes of expansions of Lucas rod sets to a rod set with a two element shape.
-Theorem 55.
+Theorem 11.3.
 For each of the following pairs $a<b$ of positive integers, the Lucas rod set $R$ expands via a finite rod set to a unique rod set with shape $\langle a,b\rangle$.
@@
 Proof.
-We need prove only existence; Theorem 52 guarantees uniqueness.
+We need prove only existence; Theorem 10.6 guarantees uniqueness.
 It suffices to find the expansions for the positive Lucas rod sets $1^{s},2^{t}]$.
-Then the invariance of the absolute values of train counts under odd sign swap 20 implies the expansions exist for ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}}^{s},2^{t}]$.
+Then the invariance of the absolute values of train counts under odd sign swap 5.4 implies the expansions exist for ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}}^{s},2^{t}]$.
 We will illustrate the proof using the running example $1^{3},2^{2}]$, which has net train counts starting at $\operatorname{F}(0)=1$
 $1,3,11,39,139,495,1763,6279,22363,79647,283667\ldots.$ (11.1)
-Every other count is a multiple of $3$, as predicted by Lemma 54.
+Every other count is a multiple of $3$, as predicted by Lemma 11.2.
 In a few cases below when the abstract proof is messy and adds no additional insight, we argue only with the example.
@@
   2. 2.
-    To find the expansion to shape $\langle a,a+2\rangle$ when $s=1$ or $a$ is even, we note that Lemma 54 implies those conditions hold just when $t\operatorname{F}(a-1)/s$ is an integer.
+    To find the expansion to shape $\langle a,a+2\rangle$ when $s=1$ or $a$ is even, we note that Lemma 11.2 implies those conditions hold just when $t\operatorname{F}(a-1)/s$ is an integer.
 Then we start with any rod set in Equation 11.2 and add $t\operatorname{F}(a-1)/s$ rod/antirod pairs of length $a$, expand the new antirods, and simplify.
@@
     To find more expansions with shape lengths that differ by $d=4$, we start with the rod set $4^{161},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{8}}^{16}]$ and repeatedly expand the smallest rods.
-Corollary 22 implies the expansions we create are also expansions of $1^{3},2^{2}]$ by finite $Q$.
+Corollary 5.6 implies the expansions we create are also expansions of $1^{3},2^{2}]$ by finite $Q$.
 Below are the first few such expansions in our example:
@@
 Every third count ($11,495,22363\ldots$) is divisible by $\operatorname{F}(2)=11$, every fourth count ($39,6279,\ldots$) by $\operatorname{F}(3)=39$, and so on.
-Definition 56.
+Definition 11.4.
 An integer sequence $\mathscr{L}$ starting at index 1 is a divisibility sequence when
 $m\mid n\implies\mathscr{L}(m)\mid\mathscr{L}(n).$
-Theorem 57.
+Theorem 11.5.
 The net train counts for a Lucas rod set $R$, reindexed so that $\mathscr{L}(n)=\operatorname{F}(n-1,R)$, form a divisibility sequence.
@@
 $(m-1)\mid(n-1)\implies\operatorname{F}(m-1)\mid\operatorname{F}(n-1).$ (11.4)
-Theorem 55 proves the existence of the expansions, which together with Equation 10.4 in the 2-shape Theorem 52 imply the needed divisibility relationships between net train counts:
+Theorem 11.3 proves the existence of the expansions, which together with Equation 10.4 in the 2-shape Theorem 10.6 imply the needed divisibility relationships between net train counts:
   1. 1.
     The expansions to rod sets with lengths $\langle a,a+2\rangle$ for odd $a$ imply the divisibilities $\operatorname{F}(1)=s\mid\operatorname{F}(a)$ for odd $a$.
-This is not a surprise, since we needed Lemma 54 to find those expansions.
+This is not a surprise, since we needed Lemma 11.2 to find those expansions.
   2. 2.
@@
 It’s somewhat less well known that the Lucas sequences, as we have defined them are also strong divisibility sequences [1][6][7].
-This result implies that the expansions we found in Theorem 55 lead to all possible two-term linear recursions involving Lucas sequences.
+This result implies that the expansions we found in Theorem 11.3 lead to all possible two-term linear recursions involving Lucas sequences.
 In particular, we have found all such Fibonacci recursions.
@@
 […]
-Definition 58.
+Definition 12.1.
 A Borwein polynomial is a polynomial all of whose coefficients are $\pm 1$. We will assume in addition that the constant term is 1.
@@
 […]
-Definition 59.
+Definition 12.2.
 Finite rod sets $S$ and $S^{\prime}$ are congruent modulo $n$ when there is a bijection $\pi$ from $S$ to $S^{\prime}$ such that for all $k\in S$, $\pi(k)\equiv k\pmod{n}$ and have the same sign.
-Theorem 60.
+Theorem 12.3.
 […]
@@
 We write down the period $6$ or period $3$ net train counts for each rod set.
-Since the counts are all $\pm 1$ or 0 Theorem 52 says the expansions correspond to a scaling that matches 1’s or 1-1s at (ba1,b1)(b-a-1,b-1) with $\alpha=1$ or one that matches 1 and 1-1 with $\alpha=-1$.
+Since the counts are all $\pm 1$ or 0 Theorem 10.6 says the expansions correspond to a scaling that matches 1’s or 1-1s at (ba1,b1)(b-a-1,b-1) with $\alpha=1$ or one that matches 1 and 1-1 with $\alpha=-1$.
 ∎
 13 $2$-Expandability when $\max R>2$
-So far we have applied Theorem 52 when $R$ contains only rods of lengths 1 and 2.
+So far we have applied Theorem 10.6 when $R$ contains only rods of lengths 1 and 2.
 We end this section (and the paper) with some examples and conjectures on $2$-expansions of rod sets $R$ with $\max R>2$.
@@
   • There is more to discover on the connections between expansions of $R$ and the structure of the net train count sequence $\operatorname{F}(n,R)$.
-Theorems 50 and 52 settle those connections for expansions to shapes of cardinality 1 and 2.
+Theorems 10.4 and 10.6 settle those connections for expansions to shapes of cardinality 1 and 2.
 What are the corresponding theorems for larger shapes?